### v1

# First moments of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions and their applications to subconvexity

Abstract.
Let $F$ be a self-dual Hecke-Maaß form for ${\rm GL}(3)$ underlying the symmetric square lift of a ${\rm GL}(2)$-newform of square-free level and trivial nebentypus. In this paper, we are interested in the first moments of the central values of $\rm GL(3)\times GL(2)$ $L$-functions and $\rm GL(2)$ $L$-functions. As a result, we obtain an estimate for the first moment for $L(1/2,F\otimes f)$ over a family, where $F$ is of the level $q^{2}$, and $f\in\mathcal{B}^{\ast}_{k}(M)$ for any primes $q,M\geq 2$ such that $(q,M)=1$. We prove the subconvex bound for $L(1/2,F\otimes f)$ involving the levels aspects simultaneously in the range $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$ for the first time. Moreover, we present a first moment estimate for $L(1/2+it,F\otimes f)$ over $K\leq k\leq 2K$, with $K$ being a large number. As an application, whenever $k^{10/9+\varepsilon}\leq|t|\leq k^{2-\varepsilon}$, we obtain a non-trivial estimate for the first moment of $L(1/2,f)L(1/2+it,F\otimes f)$ in the weight $k$ and $t$ aspects.
Key words and phrases:
Automorphic forms, first moment, Rankin-Selberg $L$-functions
2010 Mathematics Subject Classification:
11F67, 11F66, 11L05
1. Introduction
1.1. Overview
The subconvexity problem for $L$-functions is a very basic topic in number theory, which plays an important rôle in arithmetic. In history, for $L$-functions of degree one or two, this problem is completely solved. There are considerable literature on subconvexity of these $L$-functions in various aspects. For degree of three or higher degrees, this becomes very challenging, and only a limited number of results were known.
Typically, a powerful way in studying subconvexity is to appeal to the moment method. See notably Ivić [20], Conrey-Iwaniec [10], Xiaoqing Li [32] and Blomer [2] for relevant descriptions and the references therein. Let $N_{1},N_{2}\geq 1$ be two integers, and $k$ an even integer. Let $F(z)$ be a normalized Hecke-Maaß form of type $\nu=(\nu_{1},\nu_{2})$ for the congruent subgroup $\Gamma_{0}(N_{1})$ with trivial nebentypus, and $f\in\mathcal{B}^{\ast}_{k}(N_{2})$ be a Heke newform on $\text{GL}(2)$ of weight $k$ and level $q$ with trivial nebentypus; see §2.2 for definitions and backgrounds. In this paper, we will be concerned about the central values of $\text{GL}(3)\times\text{GL}(2)$ Rankin-Selberg $L$-function $L(s,F\otimes f)$, and the product of $\text{GL}(3)\times\text{GL}(2)$ and $\text{GL}(2)$ $L$-functions $L(s,F\otimes f)L(s,f)$. This is a very crucial step in understanding automorphic $L$-functions of higher degrees of six and eight.
In 2012, Xiaoqing Li [32] made the first breakthrough to break the convex barriers of $\text{GL}(3)$ $L$-functions in the $t$-aspect as well as $\text{GL}(3)\times\text{GL}(2)$ $L$-functions in the spectral aspect by investigating the first moment of $\text{GL}(3)\times\text{GL}(2)$ $L$-functions. Not long after this, Blomer [2] achieved the subconvexity for $\text{GL}(3)\times\text{GL}(2)\times\text{GL}(1)$ $L$-functions and $\text{GL}(3)\times\text{GL}(1)$ $L$-functions in their conductors aspects. With many sophisticated methods being developed in recent few years such as $\delta$-method (see, for example, [38, 39, 41]), amplification method (see, for example, [17]), deep analysis of the average involving the Kloosterman sums (see, for example, [28, 44]) as well as the integral transforms (see, for example, [39]) to name a few, the researchers were gradually attempting to solve the following well-known subconvexity conjecture:
Conjecture 1.1.
Let $N_{1},N_{2}\geq 1$ be two integers. Let $k$ be an even integer and $t\geq 1$. For any normalized ${{\rm GL}}(3)$-Maaß form $F$ of level $N_{1}$ and ${{\rm GL}}(2)$-Maaß form $f$ of level $N_{2}$ and weight $k$, there holds that $L(1/2+it,F\otimes f)\ll_{\varepsilon}(N_{1}N^{3/2}_{2}k^{3}t^{3})^{1/2-\varepsilon}$ for some absolute constant $\varepsilon>0$.
We will collate some results on this conjecture in different regards.
  • In history, Khan [26] firstly broke the convex bound in the $N_{2}$-aspect. In the case of $L(1/2,F\otimes f)\geq 0$, he exploited the amplification method to establish a conditional subconvexity
    $L(1/2,F\otimes f)\ll N^{3/4-1/2001+\varepsilon}_{2},$
    when $N_{2}$ is a prime (which should be though to be the hardest case).
  • Sharma and Sawin [43] employed a type of ‘Mass transfer’ as given by Munshi [40], showing that
    $L(1/2,F\otimes f)\ll p^{3/2-1/16+\varepsilon}$
    for any $f\in\mathcal{B}_{k}(p^{2},\chi^{2})$ with the nebentypus $\chi^{2}$ being of modulus $p^{2}$ (where $p$ is a prime). Later, Kumar, Mallesham and Singh [28] gave a better exponent of $3/2-3/20+\varepsilon$ in the depth aspect, where $f\in\mathcal{B}_{k}(p^{2r},\chi^{2})$ with the nebentypus $\chi^{2}$ being of modulus $p^{r}$, $r>3$.
  • Concerning the $t$-aspect subconvexity for $\text{GL}(3)\times\text{GL}(2)$, Munshi [41] developed the $\delta$-method to its furthest achievement by successfully establishing the following
    $L(1/2+it,F\otimes f)\ll t^{3/2-1/42+\varepsilon}.$
    Subsequently, there were in droves many works devoted to the strengthened (or the generalized) subconvexity; see, for example, [33, 19, 15]. The best exponent should be $3/2-3/20+\varepsilon$ due to Lin and Sun [33].
  • Recently, Kumar [27] firstly proved the subconvex bound in the weight $k$ aspect. By using the ‘conductor lowering mechanism’ developed by Munshi [39], he showed that
    $L(1/2+it,F\otimes f)\ll k^{3/2-1/51+\varepsilon}.$
  • Regarding the level aspect subconvexity involving the $\text{GL}(3)$ forms, it is completly out of reach. By currently developed techniques, it appears that one cannot get any power-saving to beat the subconvex barrier whatsoever. More recently, Kumar, Munshi and Singh [29] were able to prove that
    $L(1/2,F\otimes f)\ll(N_{1}N^{3/2}_{2})^{1/2+\varepsilon}\left(\frac{N^{1/4}_{1}}{N^{3/8}_{2}}+\frac{N^{1/8}_{2}}{N^{1/4}_{1}}\right),$ (1.1)
    when $N_{1},N_{2}$ are both co-prime primes. Their bound is subconvex in the range $N^{1/2+\varepsilon}_{2}<N_{1}<N^{3/2-\varepsilon}_{2}$.
1.2. Main results
In this paper, we are more concerned about the self-dual ${\rm GL}(3)\times{\rm GL}(2)$ $L$-functions, which have important implications to equidistribution problems of $L^{q}$-mass of cusp forms (as $q\rightarrow\infty$) and Heegner points on Shimura curves as well as the Quantum Unique Ergodicity (QUE) conjecture (see, for example, [4, 14, 3, 35, 34, 13, 6, 42]). For instance, the well-known Watson’s formula implies
$\|g\|_{4}^{4}\,\approx\frac{1}{p^{2}}\sum_{h\in\mathcal{B}_{2k}^{\ast}(p)}L(1/2,h)L(1/2,\text{sym}^{2}g\otimes h)$
for any ${\rm GL}(2)$-newform $g\in\mathcal{B}^{\ast}_{k}(p)$ of large fixed weight $k$ and prime level $p$. Notice that the level aspect subconvexity for ${\rm GL}(3)$ forms is still wide-open for specialists, albeit the notable successes of subconvexity for ${\rm GL}(3)$ $L$-functions in various aspects over the past decade. One of main objectives of the present paper is to address the subconvexity in the levels aspects for ${\rm GL}(3)$ and ${\rm GL}(2)$ forms simultaneously [^2] One might wander whether the method of Kumar, Munshi and Singh can be adpated to obtain subconvexity for $L(1/2,\text{sym}^{2}g\otimes f)$ with $F=\text{sym}^{2}g$ being a symmetric square lift of a ${\rm GL}(2)$-newform $g$, say. Unfortunately, this seems impracticable. The underlying reason is that the modulus $\mathcal{Q}$ in equipping with $\delta$-symbol method is usually large relative to the level of $F$; see [29, Section 2-2B]. This gives rise to a deadlock since of an inevitable gap for ${\rm GL}(3)$-Voronoĭ formulae which we will explain after a while.. To be specific, we prove the following results:
Theorem 1.2.
Let $q,M\geq 2$ be two primes satisfying that $(q,M)=1$ and $q<M^{3/8}$. Let $F$ be a Hecke-Maaß form for the congruence subgroup $\Gamma_{0}(q^{2})$ obtained as the symmetric square lift of a ${\rm GL}(2)$-newform of level $q$ with trivial nebentypus. Define the set
$\mathscr{P}=\{p,p^{2}:p\text{ is a prime },(p,qM)=1,L\leq p\leq 2L\}$
with
$\frac{q^{2+\varepsilon}}{{\sqrt{M}}}+\frac{\sqrt{M}}{q^{2-\varepsilon}}\leq L\leq\min{\left\{q^{2-\varepsilon}\sqrt{M},\frac{M^{1-\varepsilon}}{q^{2}}\right\}}.$
Then, for any $\ell\in\mathscr{P}$, we have
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f)\ll(qM)^{\varepsilon}\left(\frac{1}{\sqrt{\ell}}+q\sqrt{\frac{\ell}{M}}+\frac{q^{2}\ell}{M^{5/8}}+\frac{q\ell^{2}}{M^{5/8}}+\frac{q^{7/2}\ell^{7/4}}{M^{9/8}}\right),$ (1.2)
where the implied constant depends only on the weight $k$ and the Langlands parameters $\alpha_{i}$, $1\leq i\leq 3$, of the form $F$ as given in (2.11).
Observe that under the self-duality for $\rm GL(3)$ forms, the non-negativity for the central values of the ${\rm GL}(3)\times{\rm GL}(2)$ $L$-functions is guaranteed by Lapid’s method [30]; see also [32, Remark 3]. As an application of Theorem 1.2, we are thus able to obtain the following subconvex bound:
Corollary 1.3.
Let the forms $F,f$ and the parameters $q,M,k$ be as in Theorem 1.2. Let $\mathcal{X}\geq 1$ such that $(M^{3/2}q^{2})^{1-\delta}\leq\mathcal{X}\leq(M^{3/2}q^{2})^{1+\varepsilon}$ for arbitrarily small constants $\varepsilon,\delta>0$. Then, for any $L$ with
$\frac{q^{2+\varepsilon}}{{\sqrt{M}}}+\frac{\sqrt{M}}{q^{2-\varepsilon}}+\frac{M^{1/4+\varepsilon}}{q}\leq L\leq\min{\left\{M^{3/16-\varepsilon},\frac{M^{3/8-\varepsilon}}{q}\right\}},$ (1.3)
we have
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2+\varepsilon}\left(\frac{M^{1/4}}{qL}+\frac{\sqrt{L}}{M^{1/4}}+\frac{Lq}{M^{3/8}}+\frac{L^{2}}{M^{3/8}}+\frac{L^{7/4}q^{5/2}}{M^{7/8}}+\frac{1}{\sqrt{M}}\right).$ (1.4)
This establishes the subconvexity for any $q,M$ with $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$, and should be the first instance for subconvexity in the levels aspects for ${\rm GL}(3)\times{\rm GL}(2)$. Taking a little effort, it is not hard to attain the subconvexity bounds for any $q,M$ varying in this range $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$. For instance, when $q\asymp M^{3/16}$, we have
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2-1/30+\varepsilon}.$
In this paper, another motivation is to investigate the first moments of $L(1/2+it,F\otimes f)$ and $L(1/2,f)L(1/2+it,F\otimes f)$ in the weight $k$ and $t$ aspects over families. We continue to explore the first moment method by adding a ‘harmonic’ average of the weight $k$, with $K<k\leq 2K$. As a result, we are able to establish
Theorem 1.4.
Let $F(z)$ be a normalized Hecke-Maaß form on ${{\rm GL}}(3)$ (not necessarily self-dual). Let $k$ be an even integer. Let $K$ be a sufficiently large parameter and $K^{5/9+\varepsilon}<|t|\ll K^{2-\varepsilon}$. Let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Then, for any integer $\ell\geq 1$ satisfying that $K^{1+\varepsilon}/|t|^{3}\leq\ell\leq|t|^{3-\varepsilon}/K^{7/3}$, and any ${{\rm GL}}(2)$-newform $f$ of weight $k$ and full level, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2+it,F\otimes f)$
$\hskip 122.34692pt\ll(K|t|)^{\varepsilon}\left(\ell^{7/16}K^{27/16}|t|^{27/16}+\frac{\ell^{5/8}|t|^{23/8}}{K^{1/8}}\right).$ (1.5)
Consequently, as a by-product we can obtain a cute result as follows
Corollary 1.5.
Let the forms $F,f$, the parameters $k,K$ and the function $W$ be as in Theorem 1.4. Then, for any ${{\rm GL}}(2)$-newform $f$ of weight $k$ and full level, and $K^{10/9+\varepsilon}\leq|t|\leq K^{2-\varepsilon}$, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2+it,F\otimes f)\ll(K|t)^{\varepsilon}K^{21/8}|t|^{27/16}.$ (1.6)
This bound is already able to beat the convexity barrier in the weight $k$ and $t$ aspects, when they are varying in the range $k^{10/9+\varepsilon}\leq t\leq k^{2-\varepsilon}$ with $k$ being sufficiently large. By the convexity, the left-hand side of (1.6) is $\ll k^{3+\varepsilon}|t|^{3/2}$; while, the right-hand side shows a smaller magnitude by $\ll k^{21/8+\varepsilon}|t|^{27/16}$. Although one cannot expect the individual subconvexity bound for $L(1/2,f)L(1/2+it,F\otimes f)$, the estimate (1.4), however, reflects the effectiveness by further averaging over the parameter $k\sim K$, when it comes to investagting the subconvexity problem involving the weight aspect. We also refer the readers to the works of Luo [36] and Khan [25] for relevant heuristics. It is, on the other hand, believed that our method can be adapted to produce simultaneous non-vanishing of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions at the central point, and even subconvexity for $L(1/2,F\otimes f)$ or $L(1/2,f)L(1/2,F\otimes f)$ in the weight and the level (of $F$ or $f$) aspects. We are now working on this direction, and expect to report the progress in the future article.
The initial manoeuvre of our proofs of main results follows the path pioneered by Xiaoqing Li and Blomer, who looked into the average of first moment of $\rm GL(3)\times GL(2)$ $L$-function over a family. The strategies we pursued in the paper, however, can be understood as the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Bilinear form” in proving Theorem 1.2 and “Kuznetsov-Voronoĭ-Exponential sums” in proving Theorem 1.4, respectively. In particular, for the former, this is reminiscent of an important work of Blomer and Khan, [5], in which they investigated the twisted moments of $\text{GL}(3)\times\text{GL}(2)$ and $\text{GL}(2)$ automorphic $L$-functions, and established the following reciprocity formula
$\frac{1}{\sqrt{q}}\sum_{f\text{ level }q}L(1/2,F\otimes f)L(1/2,f)\lambda_{f}(\ell)\,\leftrightsquigarrow\,\frac{1}{\sqrt{\ell}}\sum_{f\text{ level }\ell}L(1/2,F\otimes f)L(1/2,f)\lambda_{f}(q)$ (1.7)
by the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Kuznetsov” law. Among other things, one salient point is the application of the reciprocity law, which, as the Kloosterman sum of dimension one, keeps us going back to somewhere. However, when applying the Voronoĭ to the Kloosterman sum $S(n\ell,1;c)$ of dimension two, the inverse of the variable $\ell$, that is, the exponential factor like $S(n\overline{\ell},\sigma;c)$ for some less significant parameter $\sigma$, will be occurred; this will increase the dimensions of Kloosterman sums afterwards, if one were to resort to the Voronoĭ again. In this sense, we can barely expect the reciprocity formula for Hecke-Maaß forms of higher degrees $d\geq 4$ in the levels aspects. It seems that the possible symmetric relation like (1.7) for $\text{GL}(d)\times\text{GL}(2)$ involving the levels aspects is only for the $\text{GL}(3)\times\text{GL}(2)$ case (we refer the interested readers to Jana and Nunes’s recent work [23] for the genernlization to $\rm{PGL}_{d+1}\times\rm{PGL}_{d-1}$). In contrast to the $\text{GL}(3)\times\text{GL}(2)$ case, we notice that, for the $\text{GL}(2)\times\text{GL}(1)$ case, Conrey [9] proved the conductors aspects reciprocity formula
$\frac{1}{\sqrt{p}}\sum_{\chi\bmod p}|L(1/2,\chi)|^{2}\chi(\ell)\,\leftrightsquigarrow\,\frac{1}{\sqrt{\ell}}\sum_{\chi\bmod\ell}|L(1/2,\chi)|^{2}\chi(-p).$
This, however, was further extended by Young [46] and Bettin [1] subsequently.
Let us now elucidate two of main obstacles in present paper. It is remarkable that, by the current development of the Voronoĭ formulae for ${\rm GL}(3)$ in the level aspect, there is still a gap for arbitrary non-zero integer $c\in\mathbb{Z}$ (see Lemma 2.1 below). If the additive twist colludes with the level $N$ of the form $F$, namely the parameters $c,N$ satisfy that $(N/(N,c),c)\neq 1$. For instance, one takes $N=q^{2}$ and $c=c^{\prime}q$ with $(c^{\prime},q)=1$ for any prime $q$. The ${{\rm GL}}(3)$-Voronoĭ formula in this scase looks rather tricky. In principle, Corbett’s formula [8, Theorem 1.1] can cover this, but it requires a non-trivial analysis of the $p$-adic Bessel transforms, which, however, becomes more involved; see [8] for relevant details. We also refer the readers to the work of Booker, Milinovich and Ng [7], who worked out the Voronoĭ formula in the ${{\rm GL}}(2)$-case (if the additive twist colludes with the level). As an innovation in this paper, we are able to overcome this deadlock, which makes the hybrid subconvexity for ${\rm GL}(3)$ and ${\rm GL}(2)$ in the levels aspects come into being. The key point, however, is to invoke the reciprocity law as mentioned above. Indeed, with a suitable constraint for the parameter $\ell$ we are in the unramified case, when applying the Voronoĭ for the first time; then, one would be transferred to the ramified case after an implementation of the reciprocity law. This makes it possible to equip with the Voronoĭ again.
The second obstacle is the root number. It is well-known that the explicit determination of root numbers play a focal rôle when doing averages of central values of $L$-functions over different families. In this paper, we will give a detailed description of the root number $\varepsilon(F\otimes f)$ for any ${\rm GL}(3)$ Hecke-Maaß form $F$ with arbitrary level $N$ in the framework of the adelic interpretations; see §2.3. Part of descriptions in the symmetric square lift case has already been given by Holowinsky, Munshi and Qi [16]. As it can be seen later, the root number $\varepsilon(F\otimes f)$ depends only on the parameter $k$ and the form $F$. One of its factor $i^{k}$ will have a direct influence on the ‘harmonic’ averages of the Bessel function $J_{k-1}$ in considering the moments of $L$-values in the weight $k$ aspect; see Lemmas 2.8 and 2.9.
To expose everything as clearly as possible, we have assumed that the parameter $M$ is a prime, which simplifies the argument when applying Petersson trace formula, but our method could in principle be amenable to handling the general case of square-free level with a bit more effort; see the beginning of §3.2 for relevant details.
Acknowledgements. The author is very grateful to Profs. Peter Humphries and Guangshi Lü for helpful discussions during the preparation of this work.
Notations.
Throughout the paper, $\varepsilon$ (resp. $\delta$) always denotes an arbitrarily small positive constant which may not be the same at each occurrence. $A\ll_{\varepsilon}B$ means that $|A|\ll X^{\varepsilon}|B|$, and $n\sim N$ means that $N<n\leq 2N$ for every $n\in\mathbb{Z}$. As usual, the symbols $\mathbb{R}$, $\mathbb{C}$, $\mathbb{Z}$ and $\mathbb{N}^{+}$ are respectively denote the real number field, the complex number field, the ring of integers and the ring of positive integers.
2. Prerequisites
The purpose of this section is to develop (or supply) certain preliminaries which will be assembled in the bodies of proofs of main results.
2.1. Automorphic forms
We will give a recap of some fundamental facts about cuspidal forms (see, for example, Iwaniec and Kowalski’s book [21]). Let $k\geq 2$ be an even integer and $M>0$ be a square-free integer. Let $\chi$ be a primitive character of modulus $r$ such that $M|r$, which satisfies that $\chi(-1)=(-1)^{k}$. We denote by $\mathcal{S}_{k}(M,\chi)$ the vector space of holomorphic cuspidal forms on $\Gamma_{0}(M)$ with nebentypus $\chi$ and weight $k$. For any $f\in\mathcal{S}_{k}(M,\chi)$, one then has a Fourier expansion
$f(z)=\sum_{n\geq 1}\psi_{f}(n)n^{\frac{k-1}{2}}e(nz)$
for $z\in\mathbb{H}$. Here, $e(z)$ is shorthand for $e^{2\pi iz}$ for any $z\in\mathbb{C}$, and $\mathbb{H}$ means the upper half-plane. The space $\mathcal{S}_{k}(M,\chi)$ is a finite dimensional Hilbert spaces which can be equipped with the Petersson inner products
$\langle f_{1},f_{2}\rangle=\int_{\Gamma_{0}(N)\backslash\mathbb{H}}f_{1}(z)\overline{f_{2}(z)}y^{k-2}\mathrm{d}x\mathrm{d}y.$
We recall the Hecke operators $\{T_{n}\}$ with $(n,M)=1$, which satisfy the multiplicativity relation
$T_{n}T_{m}=\sum_{d|(n,m)}\chi(d)T_{\frac{nm}{d^{2}}}.$ (2.8)
Moreover, for any $f_{1},f_{2}\in\mathcal{S}_{k}(N,\chi)$, one has
$\langle T_{n}f_{1},f_{2}\rangle=\overline{\chi(n)}\langle f_{1},T_{n}f_{2}\rangle.$
One might thus find an orthogonal basis $\mathcal{B}_{k}(M,\chi)$ of $\mathcal{S}_{k}(M,\chi)$ consisting of common eigenfunctions of all the Hecke operators $T_{n}$ with $(n,N)=1$. For each $f\in\mathcal{B}_{k}(M,\chi)$, denote by $\lambda_{f}(n)$ the $n$-th Hecke eigenvalue, which satisfies
$T_{n}f(z)=\lambda_{f}(n)f(z)$
for all $(n,M)=1$. From (2.8), one has seen that
$\psi_{f}(m)\lambda_{f}(n)=\sum_{d|(n,m)}\chi(d)\psi_{f}\left(\frac{mn}{d^{2}}\right)$
for any $m,n\geq 1$ with $(n,M)=1$. In particular, $\psi_{f}(1)\lambda_{f}(n)=\psi(n)$, if $(n,M)=1$. Therefore,
$\overline{\lambda_{f}(n)}=\overline{\chi(n)}\lambda_{f}(n),\quad\lambda_{f}(m)\lambda_{f}(n)=\sum_{d|(n,m)}\chi(d)\lambda_{f}\left(\frac{mn}{d^{2}}\right),$ (2.9)
if $(mn,M)=1.$
The Hecke eigenbasis $\mathcal{B}_{k}(M,\chi)$ also contains a subset of newforms $\mathcal{B}^{\ast}_{k}(M,\chi)$, those forms which are simultaneous eigenfunctions of all the Hecke operators $T_{n}$ for any $n\geq 1$, and normalized to have first Fourier coefficient $\psi_{f}(1)=1$. The elements of $\mathcal{B}^{\ast}_{k}(M,\chi)$ are usually called primitive forms in the sense of Atkin-Lehner; that is, each one is orthogonal to all oldforms, and is an eigenfunction of all the Hecke and Atkin-Lehner operators.
2.2. Automorphic $L$-functions and functional equations
In this part, we shall further shed some light on automorphic $L$-functions. Let $k\geq 2$ be an even integer. We assume that $N\geq 1$ is a positive integer, satisfying that $(N,M)=1$. For any newform $f\in\mathcal{B}_{k}^{\ast}(M)$, we define the attached $L$-function $L(f,s)$, for $\text{Re}(s)>1$, by the Dirichlet series $\sum_{n\geq 1}\lambda_{f}(n)n^{-s}$ which can be expressed in terms of an Euler product
$\prod_{p\nmid M}\left(1-\frac{\lambda_{f}(p)}{p^{s}}+\frac{1}{p^{2s}}\right)^{-1}\prod_{p\mid M}\left(1-\frac{\lambda_{f}(p)}{p^{s}}\right)^{-1}.$
We proceed by defining the completed $L$-function
$\Lambda(s,f)=\mathcal{Q}^{s/2}_{f}L_{\infty}(s,f)L(s,f),$
with
$L_{\infty}(s,f)=\pi^{-s}\Gamma{\left(\frac{s}{2}+\frac{k-1}{4}\right)}\Gamma{\left(\frac{s}{2}+\frac{k-1}{4}\right)}$
and the (analytic) conductor $\mathcal{Q}_{f}$ being $\mathcal{Q}_{f}\asymp M$. The completed $L$-function thus admits an analytical continuation to all $s\in\mathbb{C}$, and satisfies that
$\Lambda(s,f)=\varepsilon(f)\Lambda(1-s,f),$
where the root number $\varepsilon(f)=i^{k}\mu(M)\lambda_{f}(M)\sqrt{M}=\pm i^{k}$.
Let us now put
$\Gamma_{0}(N)=\left\{g\in\text{GL}(3,\mathbb{Z}):g\equiv\left(\begin{matrix}\ast&\ast&\ast\\ \ast&\ast&\ast\\ 0&0&\ast\end{matrix}\right)\mod N\right\}.$ (2.10)
Let $F(z)$ be a normalized Hecke-Maaß form of type $\nu=(\nu_{1},\nu_{2})$ for the congruent subgroup $\Gamma_{0}(N)$ with trivial nebentypus, which has a Fourier-Whittaker expansion with the Fourier coefficients $A_{F}(m,n)$. The Fourier coefficients of $F$ and that of its contragredient $\widetilde{F}$ are related by $A_{F}(m,n)=A_{\widetilde{F}}(n,m)$ for any $(mn,N)=1$, with $A_{F}(1,1)=1$; see, for example, [11, Chapter 6] for definition and backgrounds. The Jacquet-Langlands $L$-function is given by $L(s,F)=\sum_{n\geq 1}A_{F}(n,1)n^{-s}$ for $\text{Re}(s)>1$. Now, let
$\alpha_{1}=-\nu_{1}-2\nu_{2}+1,\quad\alpha_{1}=-\nu_{1}+\nu_{2},\quad\alpha_{1}=2\nu_{1}+\nu_{1}+1$ (2.11)
be the Langlands parameters of $F$. We define the completed $L$-function
$\Lambda(s,F)=\mathcal{Q}^{s/2}_{F}\varepsilon(F)L_{\infty}(s,F)L(s,F),$
where
$L_{\infty}(s,F)=\pi^{-3s/2}\prod_{i=1}^{3}\Gamma{\left(\frac{s+\alpha_{i}}{2}\right)},$
the (analytic) conductor $\mathcal{Q}_{F}\asymp N$, and $|\varepsilon(F)|=1$. One thus have the following functional equation
$\Lambda(s,F)=\varepsilon(F)\Lambda(1-s,\widetilde{F}).$
Next, for any $f\in\mathcal{B}^{\ast}_{k}(M)$, we turn to considering the $\text{GL}(3)\times\text{GL}(2)$ Rankin-Selberg $L$-function $L(s,F\otimes f)$, which is given by the Dirichlet series $\sum_{m,n\geq 1}A_{F}(n,m)\lambda_{f}(n)(nm^{2})^{-s}$ for $\text{Re}(s)>1$. Now, one defines the Gamma factor $L_{\infty}(s,F\otimes f)$ by
$\begin{split}L_{\infty}(s,F\otimes f)&=\pi^{-3s}\prod_{i=1}^{3}\Gamma{\left(\frac{s+{(k-1)}/{2}+\alpha_{i}}{2}\right)}\Gamma{\left(\frac{s+{(k+1)}/{2}+\alpha_{i}}{2}\right)}\end{split}$
and the (analytic) conductor $\mathcal{Q}_{F,f}\asymp N^{2}M^{3}$. The completed $L$-function defined by
$\Lambda(s,F\otimes f)=\mathcal{Q}^{s/2}_{F,f}L(s,F\otimes f)$
is thus an entire function with an analytic continuation to all $s\in\mathbb{C}$, and satisfies the function equation that
$\Lambda(s,F\otimes f)=\varepsilon(F\otimes f)\Lambda(1-s,\widetilde{F}\otimes f).$
2.3. Epsilon constant
Let the parameters $k,N,M$ be as before. Define
$k_{2,p}=\left\{g\in\text{GL}(2,\mathbb{Z}_{p}):g\equiv\left(\begin{matrix}\ast&\ast\\ 0&\ast\end{matrix}\right)\bmod p\right\}$
and
$k_{3,p}=\left\{g\in\text{GL}(3,\mathbb{Z}_{p}):g\equiv\left(\begin{matrix}\ast&\ast&\ast\\ \ast&\ast&\ast\\ 0&0&\ast\end{matrix}\right)\bmod p\right\}.$
For any given holomorphic newform $f\in L^{2}(\Gamma_{0}(M)\backslash\text{SL}(2,\mathbb{R}))$ of weight $k$, let us define a function $\varphi_{f}\in L^{2}(\text{GL}(2,\mathbb{Q})\backslash\text{GL}(2,\mathbb{A}_{\mathbb{Q}})/\text{K}(2,M))$ by
$\varphi_{f}(\gamma g_{\infty}k)=f(g_{\infty})$
for $\gamma\in\text{GL}(2,\mathbb{Q})$, $g_{\infty}\in\text{GL}(2,\mathbb{R})$ and $k\in\text{K}(2,M)$, with $\text{K}(2,M)=\text{O}(2,\mathbb{R})\times\prod_{p|M}k_{2,p}\times\prod_{p\nmid M}\text{GL}(2,\mathbb{Z}_{p})$. It can be seen that each function $\varphi_{f}$ generates the automorphic representation $\pi_{f}$ of $\text{GL}(2,\mathbb{A}_{\mathbb{Q}})$ with trivial central character. We refer the reader to, for example, [12, Sections 13.3-13.6] for relevant backgrounds.
For any given Hecke-Maaß form $F$ of level $N$ on $\text{GL}(3)$, we now let $\Pi_{F}$ be the associated automorphic representation, which belongs to the space $L^{2}(\Gamma_{0}(N)\backslash{\text{SL}}(3,\mathbb{R}))$. Here, the definition of $\Gamma_{0}(N)$ can be referred to (2.10) in §2.1. The representation $\Pi_{F}$ is generated by the smooth functions $\varphi_{\Pi_{F}}\in L^{2}(\textrm{GL}(3,\mathbb{Q})\backslash\textrm{GL}(3,\mathbb{A}_{\mathbb{Q}})/\text{K}(3,N))$, which are defined in the following way
$\varphi_{\Pi_{F}}(\gamma g_{\infty}k)=F(g_{\infty})$
for any $\gamma\in\text{GL}(3,\mathbb{Q})$, $g_{\infty}\in\text{GL}(3,\mathbb{R})$ and $k\in\text{K}(3,N)$, with $\text{K}(3,N)=\text{O}(3,\mathbb{R})\times\prod_{p|N}k_{3,p}$
×$\times\prod_{p\nmid N}\text{GL}(3,\mathbb{Z}_{p})$. It is well-known that one might factor
$\Pi_{F}=\Pi_{F,\infty}\otimes_{p<\infty}\Pi_{F,p},$
where $\Pi_{F,p}$ is an irreducible, admissible and unitary representations of $\text{GL}(3,\mathbb{Q}_{p})$ for every $p<\infty$.
We shall now develop some facts on the epsilon constants, which play a key rôle in later analysis. Observe that, for any unramified places $v=p$ with $p\nmid NM$, there already hold that $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=1$. We will focus on the archimedean place $v=\infty$ and the finite places $v=p,p|NM$ in the following.
  • To begin with, at archimedean place $p=\infty$, in the Hecke-Maaß form setting $\Pi_{F,\infty}$ is an unramified principal series representation; while, $\pi_{f,\infty}$ is a ramified discrete series representation of weight $k$, by the multiplicativity of epsilon factors we have
    $\varepsilon(\Pi_{F,\infty}\otimes\pi_{f,\infty})=\varepsilon(\pi_{f,\infty})^{3}=i^{k}.$
  • Now, for $p|M$, the local component $\Pi_{F,p}$ is an unramified principal series representation, and $\pi_{g,p}$ is the Steinberg representation, one finds
    $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=\varepsilon(\pi_{f,p})^{3}=-\lambda_{f}(p)\sqrt{p}.$
  • For $p|N$, one observes that the local component $\Pi_{F,p}$ is the Steinberg representation; while, $\pi_{g,p}$ is an unramified principal series representation with trivial central character, so that
    $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=\varepsilon(\Pi_{F,p})^{2},$
    with $|\varepsilon(\Pi_{F,p})|=1$.
We finally obtain
$\varepsilon(F\otimes f)=\prod_{p\leq\infty}\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=-i^{k}\lambda_{f}(M)\sqrt{M}\prod_{p|N}\varepsilon(\Pi_{F,p})^{2}.$ (2.12)
2.4. Voronoĭ summation formula
Let $w$ be a compactly supported smooth function on $(0,\infty)$, and $\widetilde{w}$ be its Mellin transform. For any $\rho=0,1$ and any Hecke-Maaß form $F$ of level $N$ with the Langlands parameters $\alpha_{i}$, $1\leq i\leq 3$, being as in (2.11), we define
$\gamma_{\rho}(s)=\varepsilon(F)i^{\rho}\pi^{3(s-1/2)}\prod_{j=1}^{3}\frac{\Gamma\left(\frac{1+s+\alpha_{j}+\rho}{2}\right)}{\Gamma\left(\frac{-s-\alpha_{j}+\rho}{2}\right)}$ (2.13)
and set
$\gamma_{\pm}(s)=\frac{1}{2}\left(\gamma_{0}(s)\mp i\gamma_{1}(s)\right).$
Now, we proceed to define two auxiliary quantities
$\Omega^{\pm}_{w}(x)=\frac{1}{2\pi i}\int_{(\sigma)}x^{-s}\gamma_{\pm}(s)\widetilde{w}(-s)\mathrm{d}s,$ (2.14)
where $\sigma>-1+\max\{-\text{Re}(\alpha_{1}),-\text{Re}(\alpha_{2}),-\text{Re}(\alpha_{3})\}$. We then have the following Voronoĭ summation formula for Hecke-Maaß forms on $\text{GL}(3)$ in the level aspect (see [47]).
Lemma 2.1.
Let $w(x)\in\mathcal{C}^{\infty}_{c}(0,\infty)$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Let $a,\overline{a},c\in\mathbb{Z}$ with $c\neq 0$, $(a,c)=1$, satisfying that $a\overline{a}\equiv 1\bmod c$.
  • If $(cm,N)=1$, we then have
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{a\overline{n}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sqrt{N}\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am\overline{q},\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}Nm};w\right).\end{split}$ (2.15)
  • If $N|c$, we, however, have
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{a\overline{n}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am,\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}m};w\right).\end{split}$ (2.16)
One has a good control of the asymptotic behavior of $\Omega_{\pm}$, upon expressing it in term of the oscillatory integral; see, for example, [31, Lemma 6.1].
Lemma 2.2.
For any fixed integer $\mathcal{K}\geq 1$ and $x\gg 1$, we have
$\begin{split}&\Omega_{+}(x;w)=x\int_{\mathbb{R}^{+}}w(y)\sum_{j=1}^{\mathcal{K}}\frac{1}{(xy)^{{j}/{3}}}\left[c_{j}\,e{\left(3(xy)^{1/{3}}\right)}\right.\\ &\phantom{=\;\;}\left.\hskip 163.03416pt+d_{j}\,e{\left(-3(xy)^{1/{3}}\right)}\right]\mathrm{d}y+O{\left(x^{({-\mathcal{K}+2})/{3}}\right)},\end{split}$
where $c_{j},d_{j}$ are suitable constants depending on the four parameters $\alpha_{i}$, $1\leq i\leq 3$, above. Furthermore, $\Omega_{-}(x;\omega)$ has the same expression except values of constants.
This, in turn, reduces the right-hand side of the Voronoĭ formula (2.15) essentially to
$\frac{X^{{2}/{3}}}{cN^{1/6}}\sum_{\pm,n_{1}|c}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{(n_{2}/n_{1})^{1/{3}}}\,S(a,\pm n_{2};c/n_{1})\int_{\mathbb{R}^{+}}w(y)\,e\left(\pm\frac{3\left(n_{2}n^{2}_{1}Xy\right)^{1/{3}}}{c}\right)\mathrm{d}y.$ (2.17)
The main contribution occurs when $n_{2}\ll c^{3}/n^{2}_{1}X^{1-\varepsilon}$, by repeated integration by parts, and in which case $\Omega_{\pm}\ll X^{\varepsilon}$.
2.5. A Wilton-type bound
We record the following well-known Wilton-type bound for Hecke-Maaß forms on $\text{GL}(3)$; see [37, Theorem 1.1].
Lemma 2.3.
Let $X>0$. Let $W$ be a smooth function supported on $[1/2,5/2]$ with partial derivatives satisfying $x^{i}W(x/X)\ll_{i}ZP^{i}$ for any $i\geq 0$ and some $Z,P>0$. Then, for any $\alpha\in\mathbb{R}$ we have
$\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}e(m\alpha)\,W{\left(\frac{m}{X}\right)}\ll ZPX^{1/4+\varepsilon}.$
2.6. Petersson trace formula
For any integers $m,n\geq 1$, we set
$\Delta_{k,M}(m,n)=\sum_{f\in\mathcal{B}_{k}(M)}\omega^{-1}_{f}\psi_{f}(m)\psi_{f}(n)$
and
$\Delta^{\ast}_{k,M}(m,n)=\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(m)\lambda_{f}(n),$
where the spectral weights $\omega_{f}$ are given by $\omega_{f}:=\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\langle f,f\rangle$. In particular, if $f\in\mathcal{B}^{\ast}_{k}(M)$, one has $\omega_{f}=\frac{(k-1)M}{2\pi^{2}}L(1,\text{sym}^{2}f),$ which satisfies that $k^{1-\varepsilon}M\ll\omega_{f}\ll k^{1+\varepsilon}M$ for any $\varepsilon>0$; see [22, Lemma 2.5].
Lemma 2.4.
(Petersson trace formula) We have
$\Delta_{k,M}(m,n)=\delta(m,n)+2\pi i^{-k}\sum_{c>0}\frac{S(m,n;cM)}{cM}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{cM}\right),$
where $\delta(m,n)=1$ if $m=n$, and $\delta(m,n)=0$ otherwise.
For the purpose of our application, we need an expression for $\Delta^{\ast}_{k,M}$ instead of $\Delta_{k,M}$. Suppose that $m,n$ are positive integers satisfying $(m,M)=1$, we have the following identity (see [22, Eqn. (2.51)]):
Lemma 2.5.
We have
$\Delta^{\ast}_{k,M}(m,n)=\sum_{RS=M}\frac{\mu(S)}{L\nu((m,S))}\sum_{l|S^{\infty}}l^{-1}\sum_{l^{2}_{1}|(n,\,Sl_{1})}\mu(l_{1})l_{1}\Delta_{k,R}\left(ml^{2},\frac{n}{l^{2}_{1}}\right),$
where $\nu(M)=[\Gamma(1):\Gamma(M)]=M\prod_{p|M}(1+p^{-1})$.
2.7. Bounds on bilinear forms with Kloosterman sums
The bilinear form with Kloosterman sums always plays a vital rôle in many topics such as the moments of central values of $L$-functions, subconvexity even the classic bilinear exponential sums; see, for example, [28, 24, 44] for descriptions in different regards. In this paper, we will develop the following relevant results to facilitate our use after a while.
Lemma 2.6.
Let $X,Y\geq 1$. Let $W$ be a smooth function supported on $[1/2,5/2]\times[1/2,5/2]$ with partial derivatives satisfying $x^{i}y^{j}\frac{\partial^{i}}{\partial x^{i}}\frac{\partial^{j}}{\partial y^{j}}W(x/X,y/Y)\ll_{i,j}ZZ^{i}_{1}Z^{j}_{2}$ for some $Z,Z_{1},Z_{2}>0$. Let $q\geq 1$ be an integer. For any $h\in\mathbb{Z}$, with $(h,q)=1$, we have
$\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(n,hm;q)\,W{\left(\frac{n}{X},\frac{m}{Y},\right)}\ll Z_{1}Z\sqrt{XYq}+Z_{1}ZXq^{3/4}+Z_{1}Zq.$ (2.18)
Proof.
This lemma essentially follows from the work of Kerr, Shparlinski, Wu and Xi [24] which appealed to the approach in [44, Theorem 2.1]. By partial summation, one finds that
$\begin{split}&\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(n,hm;q)W{\left(\frac{n}{X},\frac{m}{Y}\right)}\\ &\ll\int^{5X/2}_{X/2}\left|\hskip 5.0pt\sideset{}{{}^{\ast}}{\sum}\limits_{x\bmod q}\sum_{m\geq 1}\alpha_{m,t}\,e{\left(\frac{m\overline{x}}{q}\right)}\sum_{n\leq t}e{\left(\frac{nhx}{q}\right)}\right|\mathrm{d}t,\end{split}$
with $\alpha_{m,t}=\frac{\partial}{\partial t}W(t/X,m/Y)A_{F}(m,1)/\sqrt{m}$. This is compatible with the setting as presented in the beginning of the [44, Section 5]. We still might bound the inner triple sum by ±$\sum_{\pm}\sum_{1\leq i\leq I}|S^{\pm}_{i}|$, where, for $1\leq i\leq I$, $S^{\pm}_{i}=\sum_{m\leq X}\alpha_{m,t}\gamma_{x}\,e{({\pm m\overline{x}}/{q})}$; see [24, Section 6] for details. ∎
Observe that the bilinear forms we are concerned about is a smooth-weighted sum. The direct use of the Possion shows some bonus. Indeed, we alternatively have
Lemma 2.7.
Keep the notation as in Lemma 2.6. Then we have
$\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(nh,m;q)W{\left(\frac{n}{X},\frac{m}{Y}\right)}\ll Z\sqrt{XYq}+\sqrt{Z_{2}}ZXq^{3/4}\mathbf{1}_{Z_{2}q>Y}+ZX\sqrt{Y}.$ (2.19)
Proof.
By switching the orders of the summations and applying the Cauchy-Schwarz inequality followed by the Possion, one deduces that the double sum is controlled by
$\begin{split}&\ll\left(\frac{Y}{q}\sum_{n_{1},n_{2}\geq 1}\,\sum_{|\jmath|\ll Z_{2}q/Y^{1-\varepsilon}}\Phi(\jmath;n_{1},n_{2})\sum_{\alpha\bmod q}S(n_{1},\alpha;q)\overline{S(n_{2},\alpha;q)}\,e{\left(\frac{\jmath\alpha}{q}\right)}\right)^{1/2},\end{split}$
with $\Phi(\jmath;v_{1},v_{2})=\int_{\mathbb{R}^{+}}W(v_{1},t)W(v_{2},t)e(-\jmath t/q)\mathrm{d}t$ for any $v_{1},v_{2}\in\mathbb{R}^{+}$ and $\jmath\in\mathbb{R}$. It is verifiable that the zero-frequence shows a quantity $Z\sqrt{XYq}+XZ\sqrt{Y}$; while, if $Z_{2}q>Y$, the non-zero frequences, however, will contribute the quantity $\sqrt{Z_{2}}Zq^{3/4}\sqrt{Y}$ by invoking the estimate for triple sums of Kloosterman sums; see, for example, [24, Lemma 4.1]. ∎
2.8. Bessel functions
For any $x\geq 1$ and $k\in\mathbb{N}^{+}$, usually one has the crude bound for the Bessel function $J_{k}(x)$ that
$J_{k}(x)\ll{\min{\left(x^{k},x^{-1/{2}}\right)}};$ (2.20)
see, for example, [13, Proposition 8]. For any $x\in\mathbb{R}$, one also might write
$J_{k-1}(x)=e(x)\mathscr{W}^{+}_{x}(s)+e(-x)\mathscr{W}^{-}_{k}(x),$ (2.21)
with
$\mathscr{W}^{\pm}_{k}(x)=\frac{e^{\mp\left(k/2+1/{4}\right)\pi i}}{\sqrt{2\pi x}\,\Gamma\left(k+1/{2}\right)}\int_{\mathbb{R}^{+}}e^{-t}\left(x\left(1\pm\frac{it}{2x}\right)\right)^{k-1/{2}}\mathrm{d}t,$
which satisfy that
$x^{j}{\mathscr{W}^{\pm}_{k}}^{(j)}(x)\ll_{j,k}\min\left(\frac{1}{\sqrt{x}},x^{k-1}\right)$ (2.22)
for any $j\geq 0$; see, for example, [45, Sections 16.12, 17.5 & 16.3]. Now, for any $h$, a smooth bump function supported on $[1/2,5/2]$ with bounded derivatives, we define
$\widehat{h}(t)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}^{+}}\frac{h(\sqrt{u})}{\sqrt{u}}\,e^{itu}\mathrm{d}u.$
We thus have (see [22, Corollary 8.2])
Lemma 2.8.
For any $K>0$ and $x>0$, there holds that
$\sum_{k\equiv 0\bmod 2}i^{-k}h{\left(\frac{k-1}{K}\right)}J_{k-1}(x)=-\frac{K}{2\sqrt{x}}\,\text{Im}{\left(e{\left(\frac{x}{2\pi}-\frac{1}{8}\right)}\right)}\widehat{h}{\left(\frac{K^{2}}{2x}\right)}+O{\left(\frac{x}{K^{4}}\right)},$ (2.23)
where the implied constant depends only on $h$.
This shows that the oscillatory range for $J_{k-1}(x)$ is roughly $x>k^{2}$, when averaging over the weight $k$; within this range $J_{k-1}(x)$ obeys the asymptotic behavior as shown in (2.24). The function $\widehat{h}$ behaves like a Schwarz function, and satisfies that $\widehat{h}^{(j)}(x)\ll(1+|x|)^{-A}$ for any large $A$ and $j>0$ by repeated integration by parts sufficiently many times. On the other hand, if there does not exist the a factor $i^{k}$ in (2.24), one, however, has (see [22, Proposition 8.1])
Lemma 2.9.
For $\iota=0,2$ and any $K>0$ and $x>0$, there holds that
$\begin{split}4\sum_{k\equiv\iota\bmod 4}h{\left(\frac{k}{K}\right)}J_{k-1}(x)=&h{\left(\frac{k-1}{K}\right)}-i^{-\iota}\frac{K}{\sqrt{x}}\,\text{Im}{\left(e{\left(\frac{x}{2\pi}-\frac{1}{8}\right)}\right)}\widehat{h}{\left(\frac{K^{2}}{2x}\right)}\\ &+O{\left(\frac{x}{K^{3}}\left|h^{(3)}{\left(\frac{x}{K}\right)}\right|\right)}+O{\left(\frac{x}{K^{4}}\right)},\end{split}$ (2.24)
where the implied constants depend only on $h$.
3. Proofs of Theorem 1.2 and Corollary 1.3
In this section, we are ready to deduce Theorem 1.2, by which one can further show Corollary 1.3.
3.1. Proof of Theorem 1.2
Let $q$ be a prime and $M$ is also, satisfying that $(q,M)=1$. For any $f\in\mathcal{B}^{\ast}_{k}(M)$ and $F$ being a normalized $\text{GL}(3)$ Hecke-Maaß form of level $q^{2}$ with trivial nebentypus, we will consider the first moment of the central value $L(1/2,F\otimes f)$ in the levels aspects. Recall that $\mathcal{Q}_{F,f}\asymp M^{3}q^{4}$. The functional equation method shows that
$\begin{split}L(1/2,F\otimes f)=&\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}U{\left(\frac{n}{q^{2}M^{3/2}}\right)}\\ &\hskip 65.44142pt+\varepsilon(F\otimes f)\sum_{n\geq 1}\frac{A_{\widetilde{F}}(n,1)\lambda_{f}(n)}{\sqrt{n}}V{\left(\frac{n}{q^{2}M^{3/2}}\right)},\end{split}$
up to two smooth weight functions $U,V$ satisfying $U^{{j}}(x),{V}^{{j}}(x)\ll(1+|x|)^{-A}$ for any $j>0$ and sufficiently large $A>0$. Let $L\geq 1$ be a parameter to be optimized later, satisfying that
$L\leq\sqrt{M},$ (3.25)
and
$L\geq\frac{q^{2+\varepsilon}}{{\sqrt{M}}},\quad L\geq\frac{\sqrt{M}}{{q^{2-\varepsilon}}}$ (3.26)
for any $\varepsilon>0$. We now define
=$\mathscr{L}=\{p:p\text{ is prime, }L\leq p\leq 2L,(p,qM)=1\}.$
By equipping with the amplifier
$\mathfrak{A}_{f}=\sum_{\ell\in\mathscr{L}}\alpha_{\ell}\lambda_{f}(\ell),\quad\text{with }\alpha_{\ell}=\begin{cases}\lambda_{f}(\ell),&\text{ if }\ell\in\mathscr{L},\\ -1,&\text{ if }\ell=p^{2},\,p\in\mathscr{L},\\ 0,&\text{ otherwise},\end{cases}$ (3.27)
one attempts to estimate an amplified first moment
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}|\mathfrak{A}_{f}|^{2}L(1/2,F\otimes f).$
It is clear that, by the relation $\lambda^{2}_{f}(p)-\lambda_{f}(p^{2})=1$, one sees that $|\mathfrak{A}_{f}|\gg L/\log L$. If one opens the square and applies multiplicativity of Hecke eigenvalues, finally our main focus is on
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f).$ (3.28)
Note that here and thereafter, $\ell$ is either a prime or a square of a prime satisfying the coprimalities with the parameters $q$ and $M$.
To proceed, let us recall that the root number $\varepsilon(F\otimes f)=\pm\prod_{p|q}\varepsilon(\Pi_{F,p})^{2}$ according to the foregoing descriptions in §2.3. In the $+^{\prime}$ case, one finds that, the quantity in (3.28) is boiled down to the analysis of the following sum
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}U{\left(\frac{n}{\mathcal{X}}\right)},$ (3.29)
with $\mathcal{Q}^{1/2-\delta}_{F,f}\leq\mathcal{X}\leq\mathcal{Q}^{1/2+\varepsilon}_{F,f}$ for any $\delta,\varepsilon>0$. Here, $U$ is a smooth bump function supported $[1/2,5/2]$, identical to one for $x\in[1,2]$, with derivatives satisfying that $x^{j}U^{(j)}\left({x}/{\mathcal{X}}\right)\ll_{j}1$ for any $j\geq 0$.
By the Cauchy-Schwarz inequality, together with the Rankin-Selberg bounds for the second moments of Fourier coefficients, the trivial estimate for the double sum above is $\ll_{\varepsilon}\mathcal{X}^{1/2+\varepsilon}$. This is far away from the bound $\ll\mathcal{X}^{1/2+\varepsilon}/M$ leading to the convexity. The task is thus reduced to estimating (3.29) further. Our strategy next is the Petersson trace formula. To set up for our application of this, we trivially write the expression (3.29) above as
$\sum_{\begin{subarray}{c}n\geq 1\\ (n,M)=1\end{subarray}}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta^{\ast}_{k,M}(n,\ell)U{\left(\frac{n}{\mathcal{X}}\right)}+\sum_{\begin{subarray}{c}n\geq 1\\ M|n\end{subarray}}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta^{\ast}_{k,M}(n,\ell)U{\left(\frac{n}{\mathcal{X}}\right)}.$ (3.30)
It is straightforward to see that, by the fact $|\lambda_{f}(M)|\ll M^{-1/2}$ that the second term is bounded by $\ll_{\varepsilon}\sqrt{\mathcal{X}}/M^{3/2}$. For $(n\ell,M)=1$, notice that, by Lemma 2.6 one has the following relation
$\Delta^{\ast}_{k,M}(n,\ell)=\Delta_{k,M}(n,\ell)-\frac{1}{M}\sum_{\rho|M^{\infty}}\frac{1}{\rho}\Delta_{k,1}{\left(n,\ell\rho^{2}\right)},$
so that one may plug the second term on the right-hand side above into (3.30) showing that
$\frac{1}{M}\sum_{\begin{subarray}{c}\rho|M^{\infty}\\ \rho>1\end{subarray}}\frac{1}{\rho}\sum_{\begin{subarray}{c}n\geq 1\\ (n,M)=1\end{subarray}}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}\Delta_{k,1}{\left(n,\ell\rho^{2}\right)}U{\left(\frac{n}{\mathcal{X}}\right)}\ll_{\varepsilon}\frac{\sqrt{\mathcal{X}}}{M^{2-\varepsilon}}.$
This reduces (3.29) to
$\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\left\{\Delta_{k,M}(n,\ell)-\frac{1}{M}\Delta_{k,1}(n,\ell)\right\}U{\left(\frac{n}{X}\right)}$ (3.31)
plus an acceptable error $O_{\varepsilon}{(\mathcal{X}^{1/2+\varepsilon}/M^{3/2})}$. Here, the first term in the braces should be viewed as an average over all forms of level $M$. The second, as an average over the old forms, however, will not contribute significantly compared with the newforms. Indeed, in this case one finds that
$\begin{split}&\frac{1}{M}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta_{k,q}(n,1)U{\left(\frac{n}{X}\right)}=\frac{1}{M}\sum_{f^{\star}\in\mathcal{B}^{\ast}_{k}(1)}\omega^{-1}_{f^{\star}}\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f^{\star}}(n)}{\sqrt{n}}U{\left(\frac{n}{X}\right)},\end{split}$
which is $\ll q^{1+\varepsilon}/M$ by the convexity bound $O(q^{1+\varepsilon})$ for the Rankin-Selberg $L$-function $L(s,F\otimes f^{\star})$. Our job thus reduces to
$\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta_{k,M}(n,1)+O{\left(\frac{\sqrt{\mathcal{X}}}{M^{3/2-\varepsilon}}+\frac{q^{1+\varepsilon}}{M}\right)}.$ (3.32)
Now, we are ready to apply the Petersson trace formula (Lemma 2.4) to $\Delta_{k,M}$; this produces the diagonal and off-diagonal terms of the form $\mathcal{O}(M,q)+2\pi i^{k}\mathcal{S}(M,q)$, where
$\mathcal{O}(M,q)=\sum_{n=1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}\delta(n,\ell)U{\left(\frac{n}{\mathcal{X}}\right)}$ (3.33)
which is $O(\ell^{-1/2+\varepsilon})$, and
$\mathcal{S}(M,q)=\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}U{\left(\frac{n}{\mathcal{X}}\right)}\sum_{c\geq 1}\frac{S(n,\ell;cM)}{cM}J_{k-1}\left(\frac{4\pi\sqrt{n\ell}}{cM}\right).$ (3.34)
Noticing that we are taking $k$ sufficiently large, we see that the sum over $c$ can be truncated at $c\leq\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$; otherwise, the contribution in the complementary range is negligibly small (i.e., $O_{A}(\mathcal{X}^{-A})$ for any sufficiently large $A$), with the help of the bound (2.20). By using the dyadic subdivision, we further consider the multiple sum
$\sup_{C\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c\sim C}\frac{S(n,\ell;cM)}{cM}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.35)
with ${U}_{\ell}\left(y\right)=U(y)J_{k-1}\left({4\pi\sqrt{y\mathcal{X}\ell}}/{CM}\right)$ for any $y\in\mathbb{R}^{+}$. According to the properties of $J$-Bessel functions in §2.1, it follows that
$y^{i}\ell^{j}\frac{\partial^{i}}{\partial y^{i}}\frac{\partial^{j}}{\partial\ell^{j}}U_{\ell}(y)\ll\left(\frac{\sqrt{\mathcal{X}L}}{CM}\right)^{i-1/2}$ (3.36)
for any $i,j\geq 0$.
Next, we prepare to apply the Voronoĭ summation formula, Lemma 2.1. Before that, one needs to detect the coprimality between $\ell$ and the modulus of the twist $cM$. Recall $\ell$ is a prime or a square of a prime. To begin with, when $\ell$ is prime, the coprimality $(c,\ell)=1$ is ascertained by the first constraint of (3.26), which forces that $c<\ell$. Now, we turn to considering that the case of $\ell$ being a square of a prime. Observe that in the (possibly occurred) case where $\ell=p^{2},c=pc^{\prime}$, with $p\asymp L$, by (2.20), the quantity $\mathcal{S}(M,q)$ is negligible small, unless that $\sqrt{\mathcal{X}}\gg M^{1-\varepsilon}$; if so, one finds $S(n,p^{2};pc^{\prime}M)=-S(n,1;c^{\prime}M)$ only when $(c^{\prime},p)=1$, and zero otherwise. Our object (3.35) is degenerated into
$\frac{1}{L}\sup_{C^{\prime}\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{\begin{subarray}{c}c^{\prime}\sim C^{\prime}\end{subarray}}\frac{S(n,1;c^{\prime}M)}{c^{\prime}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right).$ (3.37)
It thus suffices to deal with (3.35) with $(c,\ell)=1$ in what follows; by a similar argument the possibly-existing quantity (3.37) will provide a relative less significant contribution. Now, using (2.15), one reduces (3.35) essentially to
$\begin{split}\sup_{C\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\frac{CMq}{\sqrt{\mathcal{X}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{mr^{2}}\,e\left(\mp\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)\,\Omega_{\pm}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}};{U}_{\ell}}\right).\end{split}$ (3.38)
Here, we have applied the estimate involving Ramanujan sum for each modulus $c$ that $S(n,0;c)=\sum_{ab=c}\mu(a)\sum_{\beta\bmod c}e{\left({\beta n}/{c}\right)}$. One should notice that the additive twist does not collude with the level anymore, namely, there holds that $(cM,q)=1$; this, however, is guaranteed by the second one of pre-conditions in (3.25) which confines that $c<q$ [^4] Indeed, on account of the final restrictions for the parameter $L$ (see (3.54) below), the pre-condition, here, seems dispensable (as far as the establishment of effective ranges for subconvexity is concerned).. By the asymptotic formula Lemma 2.2 for the integral transform $\Omega_{\pm}$, we are further reduced into the following superiorem problem that
$\begin{split}\mathcal{S}(M,q)\ll\sup_{1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\Psi(C;\ell,M,q),\end{split}$ (3.39)
with
$\Psi(C;\ell,M,q)=\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\,\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}r^{2/3}}\,e\left(\mp\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)\,\widetilde{{U}_{\ell}}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}}}\right).$
Here, $\widetilde{{U}_{\ell}}(x)=\int_{\mathbb{R}^{+}}U_{\ell}(y)e(\pm 3(xy)^{1/3})\mathrm{d}y$ for any $x\in\mathbb{R}^{+}$. By the feature (2.21), it can be expressed roughly as
$\left(\frac{\sqrt{\mathcal{X}\ell}}{CM}\right)^{-1/2}\int_{\mathbb{R}^{+}}U(y)F^{+}_{k}{\left(\frac{4\pi\sqrt{y\mathcal{X}\ell}}{CM}\right)}\,e{\left(\frac{4\pi\sqrt{y\mathcal{X}\ell}}{CM}\pm 3(xy)^{1/3}\right)}\mathrm{d}y$ (3.40)
for any $x\in\mathbb{R}^{+}$, where $F^{+}_{k}(t)=\sqrt{2\pi t}\,\mathscr{W}^{+}_{k}(t)$ for any $t\in\mathbb{R}$, satisfying that $t^{j}{F^{+}}^{(j)}_{k}(t)\ll_{j,k}1$ for any $t\gg 1$ by (2.22).
3.1.1. Non-oscillatory case
In this subsection, we would firstly proceed our argument in the situation where the weight $U_{\ell}(y)$ is of the phase roughly “1”. In other words, we start by working with $C$ localizing at $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$ in (3.39). Our target thus simplifies into
$\begin{split}\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}r^{2/3}}\,e\left(\mp\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)\,\widetilde{{U}_{\ell}}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}}}\right).\end{split}$
One finds that $\widetilde{U}^{(j)}_{\ell}\ll X^{\varepsilon}$ for any $j>0$, by repeated integration by parts many times. It thus follows that one may truncate the variable $m$ at
$m\ll\mathcal{X}^{\varepsilon}+q^{2}(CM)^{3}/\mathcal{X}^{1-\varepsilon}\ll q^{2}\sqrt{\mathcal{X}}\ell^{3/2+\varepsilon}$
at the cost of a negligible error. Our next strategy is to make use of the reciprocity law, followed by employing the GL(3)-Voronoĭ; this would lead to a dual form of (3.35). To be precise, one writes
$e\left(\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)=e\left(-\frac{mr^{2}\overline{cM}}{\ell q^{2}}\right)e\left(\frac{mr^{2}}{cM\ell q^{2}}\right).$
Observe that the harmonic $e\left({mr^{2}}/{cM\ell q^{2}}\right)$ is automatically flat. Indeed, this is not surprising, since the variables in the Kloosterman sums always lie in the “Linnik range”. It thus suffices to evaluate
$\begin{split}\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}r^{2/3}}\,e\left(\pm\frac{mr^{2}\overline{cM}}{\ell q^{2}}\right)\,\widetilde{{U}_{\ell}}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}}}\right).\end{split}$ (3.41)
For any given $\mathscr{M}\geq 1$, we introduce a smooth function $\eta_{\mathscr{M}}(m)$ such that $\sum_{\mathscr{M}\geq 1}\eta_{\mathscr{M}}(m)=1$, where $\eta_{\mathscr{M}}(m)$ is supported on $m\sim\mathscr{M}$, with $\mathscr{M}$ running through the powers of 2 independently, and satisfies that $\mathscr{M}^{j}\eta_{\mathscr{M}}^{(j)}(n)\ll\mathcal{X}^{\varepsilon}$ for any $j\geq 0$. Thus, if now one defines
${\mathcal{W}_{\pm,r,\ell}}{\left({\rho},\nu\right)}=\eta_{\mathscr{M}}(\rho\mathscr{M})\,\widetilde{U}_{\ell}\left({\frac{\rho\mathscr{M}r^{2}\mathcal{X}}{(\nu CM)^{3}q^{2}}}\right)$
for any ${\rho},\nu\in\mathbb{R}^{+}$, then we can recast (3.41) as
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\sum_{r|cM}\frac{1}{r^{2/3}}\,\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\left|\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\right|,$ (3.42)
where
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)=\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}}\,e\left(\pm\frac{mr^{2}\overline{cM}}{\ell q^{2}}\right)\,{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{m}{\mathscr{M}};\frac{c}{C}\right)}$ (3.43)
We will merely consider the $+^{\prime}$ case; the case $-^{\prime}$ can be handled in an entirely analogously manner. Note that the coprimality relation $(r^{2}\overline{cM},\ell q^{2})=1$ has already held. We now resort to the Voronoĭ formula in the ramified case, (2.16), which further converts the double sum $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ essentially into
$\begin{split}&\frac{\mathscr{M}^{1/3}}{\ell q^{2}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{(n/n^{\prime})^{1/3}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\Omega_{\pm}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};{\mathcal{W}_{\pm,r,\ell}}\right)}.\end{split}$
This, however, is controlled by
$\begin{split}&\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)},\end{split}$ (3.44)
where the integral $\widetilde{\mathcal{W}_{\pm,r,\ell}}$ is given by
$\widetilde{\mathcal{W}_{\pm,r,\ell}}(v;\iota)=\int_{\mathbb{R}^{+}}{\mathcal{W}_{\pm,r,\ell}}{\left(y;\iota\right)}\,e{\left(\pm 3(vy)^{1/3}\right)}\mathrm{d}y$
for any $v,\iota\in\mathbb{R}^{+}$. Appealing to (3.36), one temporarily sees that $\frac{{\partial^{i}}{\partial^{j}}}{{\partial v^{i}}{\partial\iota^{j}}}\,\widetilde{\mathcal{W}_{\pm,r,\ell}}(v;\iota)\ll_{i,j}\mathcal{X}^{\varepsilon}$ for any $i,j,\varepsilon>0$.
Observe that, unlike other scenarios where one only cares about what happens while the parameter $\mathscr{M}$ varying around its height by directly adding a smooth weight after the Cauchy-Schwarz (see, for example, [41, 18]), we need to optimize $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ for any $1\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$. Next, we will proceed the argument by distinguishing two different situations.
When $\mathscr{M}$ is small. We first consider the case where $\mathscr{M}$ is small such that $1\leq\mathscr{M}\leq\mathscr{M}^{\ast}$, say, with $\mathscr{M}^{\ast}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$. Recall (3.43). We then have
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{5/7}C^{1+\varepsilon}.$ (3.45)
Here, we have used the estimate $|A_{F}(n,1)|\ll n^{5/7+\varepsilon}$ which follows form the Luo-Rudnick-
Sarnak bound for the generalized Ramanujan conjecture; see, for example, [11, Therem 12.5.1].
When $\mathscr{M}$ is large. Next, let us turn to the case where $\mathscr{M}$ is suitably large, with $\mathscr{M}^{\ast}<\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$. We are in a position to rely on the estimates for bilinear Kloosterman sums securing further cancellations. For the sake of the clarity of presentation, one might merely take care of the salient case where $n^{\prime}=1$ (this will be responsible for the dominated contribution); while, for the degenerate case where $n^{\prime}$ is suitably large, our analysis works as well, and the bound that we obtain is dominated by the former (as one will expect). It turns out that the quantity we are really concerned about is the following
$\begin{split}&\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell}}\sum_{c\sim C}\sum_{n\geq 1}\frac{A_{F}(1,n)}{\sqrt{n}}\,S(\overline{r^{3}}Mc,\pm n;\ell q^{2})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)}.\end{split}$
It is necessary that ${\ell^{3}q^{6}}\gg{(CM)^{3}q^{2}}/\mathcal{X}^{1-\varepsilon}$ for $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$, so that the sum over $n$ is not vacuous. Now, an application of Lemma 2.6, with $Z=\mathcal{X}^{\varepsilon}$ and $Z_{1}=Z_{2}=1$, reveals that the display above is dominated in the magnitude by
$\begin{split}&\ll\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell}}\left\{(\ell q^{2})^{2}\sqrt{\frac{Cr}{\mathscr{M}}}+(\ell q^{2})^{3/4}C+\ell q^{2}\right\}\\ &\ll\frac{\sqrt{Cr}(\ell q^{2})^{3/2}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].\end{split}$
Combining with (3.45), it follows that
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{5/7}C^{1+\varepsilon}+\frac{\sqrt{Cr}(\ell q^{2})^{3/2+\varepsilon}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
Upon choosing
$\mathscr{M}^{\ast}=\frac{(\ell q^{2})^{3/2}}{\sqrt{C}r^{3/14}},$
it turns out that
$\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll\ell q^{2}C^{2/3+\varepsilon}r^{4/7}+\frac{\sqrt{CM}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
Inserting this expression into (3.44), we obtain
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\,\sum_{r|cM}\frac{1}{r^{2/3}}\left(\ell q^{2}C^{2/3}r^{4/7}+\frac{\sqrt{CM}q^{1/3+\varepsilon}}{(r\mathcal{X})^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right]\right).$
This implies that the portion for $r$ being large is less dominated by the focal case of $r=1$. We thus have
Lemma 3.1.
For any $\ell\sim L\geq 1$ satisfying that
$L\geq\frac{\sqrt{M}}{q^{2-\varepsilon}},\quad{L^{3}q^{6}}\geq\frac{(CM)^{3}q^{2}}{\mathcal{X}^{1-\varepsilon}},\quad\frac{(Lq^{2})^{3/2}}{\sqrt{C}}\leq\frac{q^{2}(CM)^{3}}{\mathcal{X}^{1+\varepsilon}},$ (3.46)
with $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$, we have
$\Psi(C;\ell,M,q)\ll\frac{\ell q^{5/3}\mathcal{X}^{1/6+\varepsilon}}{M}+q^{1+\varepsilon}\sqrt{\frac{\ell}{M}}+\frac{\ell^{2}\mathcal{X}^{1/4+\varepsilon}\sqrt{q}}{M}.$ (3.47)
One notices that the last two conditions in (3.46) requires that
$L\geq\frac{\mathcal{X}^{1/3+\varepsilon}}{q^{8/3}}\quad\text{and}\quad L^{5/6}\geq\frac{\mathcal{X}^{1/6}}{q^{2}M^{1/3-\varepsilon}},\quad L^{3/2}\leq\frac{M^{3}}{q\mathcal{X}^{1+\varepsilon}},$ (3.48)
respectively.
3.1.2. Oscillatory case
Next, in this part we will focus on the situation where $C$ is suitably small such that $C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$ in (3.39). Recall (3.40). In this case, the phase of the weight $U_{\ell}(y)$ will be roughly $\sqrt{\mathcal{X}\ell}/CM\gg\mathcal{X}^{\varepsilon}$. Set $A:=4\pi\sqrt{\mathcal{X}\ell}/CM$, and put $\mathscr{J}(y)=$
$U(y)F_{k}^{+}(A\sqrt{y})$ for any $y\in\mathbb{R}^{+}$. To begin with, let us take a glance at the integral
$\int_{\mathbb{R}^{+}}\mathscr{J}(y)e(A\sqrt{y}-3(yx)^{1/3})\mathrm{d}y.$
Notice that $\mathscr{J}$ is compactly supported on $[1/2,5/2]$, and satisfies that $\mathscr{J}^{(j)}\ll_{j,k}1$ for any $j>0$. On finds this integral is negligibly small, unless that $A\asymp x^{1/3}$, and, in which case it is essentially given by
$x^{-1/6}e(-4x/A^{2})\mathscr{J}{\left((2x^{1/3}/A)^{6}\right)}.$
This leads us to the multiple sum
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,1)}}{\sqrt{mr^{2}}}\,e\left(\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)e{\left(-\frac{mr^{2}}{2\pi^{2}cM\ell q^{2}}\right)}\mathscr{J}{\left({\frac{mr^{2}}{\mathcal{M}}}\right)},$
with $\mathcal{M}\asymp\mathcal{X}^{1/2+\varepsilon}\ell^{3/2}q^{2}$. Based on the analysis in the paragraphs above, it suffices to concentrate on $r=1$ case which, however, is though to be responsible for the dominated contribution. In other words, we are simplified into
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{c\sim C}\sum_{\begin{subarray}{c}m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,1)}}{\sqrt{m}}\,e\left(\frac{m\overline{\ell q^{2}}}{cM}\right)e{\left(-\frac{m}{2\pi^{2}cM\ell q^{2}}\right)}\mathscr{J}{\left({\frac{m}{\mathcal{M}}}\right)}.$
Recall that one has the coprimality relation $(cM,\ell q)=1$. Now, if one attempts to apply the reciprocity law and the Voronoĭ, we finally arrive at
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{n^{\prime}|\ell q^{2}}\,\sum_{c\sim C}\sum_{n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(cM,\pm n;\ell q^{2}/n^{\prime})\,\widetilde{\mathscr{J}_{\pm,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathcal{M}}{\ell^{3}q^{6}};\frac{c}{C}\right)}$ (3.49)
akin to (3.44), where the integral transform is given by
$\widetilde{\mathscr{J}_{\pm,\ell}}{\left(x;\nu\right)}=\int_{\mathbb{R}^{+}}\mathscr{J}{\left(y\right)}\,e{\left(\frac{\rho\mathcal{M}y}{\nu CM\ell q^{2}}\pm 3(yx)^{1/3}\right)}\mathrm{d}y$
for some non-zero constant $\rho\in\mathbb{R}$. Observe that the phase is till roughly $\sqrt{\mathcal{X}\ell}/CM$, and the integral $\widetilde{\mathscr{J}_{\pm,\ell}}$ is negligibly small, unless that $\mathcal{M}/(CM\ell q^{2})\asymp x^{1/3}$. This shows that the sum over $n$ is such that
$n{n^{\prime}}^{2}\asymp\mathscr{N}:=\mathcal{M}^{2}/(CM)^{3}\ll\mathcal{X}^{1+\varepsilon}\ell^{3}q^{4}/(CM)^{3}.$
To proceed, one sees that it is necessary that ${\mathcal{X}\ell^{3}q^{4}}/{(CM)^{3}}\gg 1$ for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$; this implies that $\ell\geq\sqrt{M}/q^{2-\varepsilon}$ which is already ensured by (3.26). Returning to (3.49), there are two options for $n^{\prime}$, that is, $n^{\prime}=1$ and $n^{\prime}|\ell,n^{\prime}>1$ ($n^{\prime}=q,q^{2}$ cannot occur since of that $q^{2}>\mathcal{X}^{1+100\varepsilon}\ell^{3}q^{4}/M^{3}$ by the final choice of $\ell$ as presented in (3.54)). We shall merely consider the case of $n^{\prime}=1$. The same method can be adapted to argue another type, the contribution of which is relatively small. As of now, it is clear that
$x^{i}\nu^{j}\frac{\partial^{i}}{\partial x^{i}}\frac{\partial^{j}}{\partial\nu^{j}}\widetilde{\mathscr{J}_{\pm,\ell}}{\left(x;\nu\right)}\ll(\sqrt{\mathcal{X}\ell}/CM)^{i+j}$
for any $i,j>0$. We now apply Lemma 2.7, with $Z=1,Z_{1}=Z_{2}=\sqrt{\mathcal{X}\ell}/CM$, finding that the expression in (3.49) is dominated by
$\ll\frac{1}{(\mathcal{X}\ell)^{1/4-\varepsilon}}\left\{\sqrt{\mathscr{N}\ell q^{2}c}+\left(\frac{\sqrt{\mathcal{X}\ell}}{CM}\right)^{1/2}(\ell q^{2})^{3/4}c+c\sqrt{\mathscr{N}}\right\}$
$\ll\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{7/4}q^{3}}{CM^{3/2}}+\frac{\sqrt{C}(\ell q^{2})^{3/4+\varepsilon}}{\sqrt{M}}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{5/4}q^{2}}{\sqrt{C}M^{3/2}}.$
This is controlled by
$\ll\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{5/4}q^{2}}{M^{3/2}}$
for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$. We thus deduce
Lemma 3.2.
For any $\ell\sim L,C\ll\sqrt{\mathcal{X}L}/M^{1+\varepsilon}$ with $L$ satisfying that (3.26), we obtain
$\Psi(C;\ell,M,q)\ll\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{5/4}q^{2}}{M^{3/2}}.$ (3.50)
Altogether, the various bounds in (3.32), (3.39), (3.47) and (3.50) above will result in
Proposition 3.3.
For any $\ell\sim L$ with $L$ satisfying that the constraints (3.25), (3.26) and (3.48), we have
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f)\ll\mathcal{X}^{\varepsilon}\left(\frac{1}{\sqrt{\ell}}+\frac{\mathcal{X}^{1/6}\ell q^{5/3}}{M}+q\sqrt{\frac{\ell}{M}}+\frac{\ell^{2}\mathcal{X}^{1/4}\sqrt{q}}{M}\right.$
$\phantom{=\;\;}\left.\hskip 61.45805pt+\frac{\mathcal{X}^{1/4}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4}\ell^{5/4}q^{2}}{M^{3/2}}+\frac{\sqrt{\mathcal{X}}}{M^{3/2}}+\frac{q}{M}\right).$
This shows (1.2) immediately after a little simplification, as required.
3.2. Proof of Corollary 1.3
In this section, we shall turn to the proof of Corollary 1.3. Indeed, as a direct consequence of Proposition 3.3, by summing over $\ell_{1}$ and $\ell_{2}$ in the amplifier
$|\mathfrak{A}_{f}|^{2}=\sum_{\ell_{1},\ell_{2}\in\mathscr{L}}\alpha_{\ell_{1}}\overline{\alpha_{\ell_{2}}}\lambda_{f}(\ell_{1})\overline{\lambda_{f}(\ell_{2})},$
it turns out that
$L(1/2,F\otimes f)\ll\mathcal{X}^{\varepsilon}\left(\frac{M}{L}+{Lq^{5/3}\mathcal{X}^{1/6}}+q\sqrt{{LM}}+{L^{2}\mathcal{X}^{1/4}}\sqrt{q}\right.$
$\phantom{=\;\;}\left.\hskip 114.94893pt+\frac{\mathcal{X}^{1/4}\ell^{7/4}q^{3}}{\sqrt{M}}+{\mathcal{X}^{1/4}\ell q^{3/2}}+\frac{\mathcal{X}^{1/4}\ell^{5/4}q^{2}}{\sqrt{M}}+\sqrt{\frac{\mathcal{X}}{M}}+q\right).$ (3.51)
To beat the convexity barrier, the parameter $L$ on the first line on the right-hand side of (3.2) satisfies that
$\begin{split}&L\geq\frac{M^{1+\varepsilon}}{\sqrt{\mathcal{X}}},\quad L\leq\frac{\mathcal{X}^{1/3-\varepsilon}}{q^{5/3}}\\ &L\leq\frac{\mathcal{X}^{1-\varepsilon}}{q^{2}M},\quad L\leq\left(\frac{\mathcal{X}}{q^{2}}\right)^{1/8-\varepsilon};\end{split}$ (3.52)
while, the second line requires that
$\begin{split}&L\leq\frac{\mathcal{X}^{1/7-\varepsilon}M^{2/7}}{q^{12/7}},\quad L\leq\left(\frac{\mathcal{X}}{q^{6}}\right)^{1/4-\varepsilon}\quad L\leq\frac{\mathcal{X}^{1/5-\varepsilon}M^{2/5}}{q^{8/5}}.\end{split}$ (3.53)
Recall that $(M^{3/2}q^{2})^{1-2\delta}\leq\mathcal{X}\leq(M^{3/2}q^{2})^{1+2\varepsilon}$ for any $\delta,\varepsilon>0$. Collecting all the inequalities (3.25), (3.26), (3.48), (3.52) and (3.53), we infer that the effective range for $L$ should be
$\frac{q^{2+\varepsilon}}{{\sqrt{M}}}+\frac{\sqrt{M}}{q^{2-\varepsilon}}+\frac{M^{1/4+\varepsilon}}{q}+1\leq L\leq\min{\left\{M^{3/16-\varepsilon},\,\frac{\sqrt{M}}{q^{10/7+\varepsilon}},\,\frac{M^{3/8-\varepsilon}}{q}\right\}}.$ (3.54)
This yields the range $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$, and thus gives the constraint as shown in (1.3). The assertion of Corollary 1.3 now follows by combining with (3.2).
4. Proofs of Theorem 1.4 and Corollary 1.5
In this section, we shall be dedicated to the proofs of Theorem 1.4 and Corollary 1.5.
4.1. Proof of Theorem 1.4
We shall consider the average of the first moments of $\rm\text{GL}(3)\times GL(2)$ and $\text{GL}(2)$ $L$-functions in a family with a ‘harmonic’ average of $k\sim K$. To be precise, let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Let $f$ be a $\text{GL}(2)$-newform of weight $k\in 2\mathbb{Z}$ and full level. Let $\ell\geq 1$ be an integer. Let $K$ be a sufficiently large parameter, and $K^{100\varepsilon}\ll|t|\ll K^{2-100\varepsilon}$. We attempt to obtain an upper-bound for
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2+it,F\otimes f).$ (4.55)
As described in §2.3, the root number $\varepsilon(F\otimes f)$ is dependent on the weight $k$; thus, by the function equation method, we need to turn to seeking an effective control over the following two sums
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{n^{1/2+it}}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.56)
and
$\sum_{k\equiv 0\bmod 2}i^{k}\,W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{n^{1/2+it}}V{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.57)
where $(|t|k)^{3-\delta}\leq\mathcal{Y}\leq(|t|k)^{3-\varepsilon}$ for arbitrarily small constants $\delta,\varepsilon>0$, and $V^{\star},V$ are two smooth bump functions supported $[1/2,5/2]$, identical to one for $x\in[1,2]$, with derivatives satisfying ${V^{\star}}^{(j)},V^{(j)}\ll_{j}1$ for any $j\geq 0$. We shall merely handle the second sum (4.57) in the following. The same analysis works for the first sum (4.56). The main difference is that, for the first sum, the Lemma 2.8 will be used, which, however, produces a possibly weaker bound since of more error terms thereof; for the second one, we will apply Lemma 2.9 instead after a while.
By equipping with Petersoon formula, Lemma 2.4, we split the expression in (4.57) into the diagonal term and the off-diagonal term. The diagonal term is roughly $\ll K^{1+\varepsilon}/\sqrt{\ell}$; while, for the off-diagonal term, it is presented as the following form
$2\pi\sum_{c\geq 1}\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{n\geq 1}\frac{A_{F}(n,1)}{n^{1/2+it}}\frac{S(n,\ell;c)}{c}J_{k-1}{\left(\frac{4\pi\sqrt{n\ell}}{c}\right)}V{\left(\frac{n}{\mathcal{Y}}\right)}.$
Now, one applies Lemma 2.9 as mentioned above, which yields two terms
$\mathcal{T}(K,t)=\sum_{c\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\,\sum_{n\geq 1}\frac{A_{F}(n,1)}{n^{1/2+it}}\frac{S(n,\ell;c)}{c}\,e{\left(\frac{4\pi\sqrt{n\ell}}{c}\right)}V{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.58)
and
$\widehat{\mathcal{T}}(K,t)=\sum_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll c\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\,\sum_{n\geq 1}\,\frac{A_{F}(n,1)}{n^{1/2+it}}\frac{S(n,\ell;c)}{c}W{\left(\frac{\sqrt{n\ell}}{cK}\right)}V{\left(\frac{n}{\mathcal{Y}}\right)}$ (4.59)
plus two error terms $\text{Err}^{\flat}(K,t)$ and $\text{Err}^{\natural}K,t)$, say, which can be estimated as
$\begin{split}\text{Err}^{\flat}(K,t)&\ll\frac{1}{K^{2-\varepsilon}}\sum_{n\sim\mathcal{Y}}\frac{|A_{F}(n,1)|}{\sqrt{n}}\sum_{c\sim\sqrt{\mathcal{Y}\ell}/K}\frac{|S(n,l;c)|}{c}\ll\frac{X^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}},\end{split}$ (4.60)
and
$\begin{split}\text{Err}^{\natural}(K,t)&\ll\frac{\sqrt{\ell}}{K^{4-\varepsilon}}\sum_{n\sim\mathcal{Y}}{|A_{F}(n,1)|}\sum_{c\ll\sqrt{\ell}\mathcal{Y}^{1/2+\varepsilon}}\frac{|S(n,l;c)|}{c^{2}}\ll\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}},\end{split}$ (4.61)
respectively. Here, $\text{Err}^{\flat}(K,t)$ comes from the first error term in (2.24); while, $\text{Err}^{\natural}(K,t)$ is for the second one. We have applied bound (2.20) in estimating $\text{Err}^{\natural}(K,t)$, when $k$ is sufficiently large.
In the following paragraphs, we will be devoted to the estimates of $\mathcal{T}$ and $\widehat{\mathcal{T}}$ given as in (4.58) and (4.59), respectively.
4.1.1. Bounding $\mathcal{T}$
We first take care of $\mathcal{T}$. By the dyadic subdivision, what we need to be concentrated on actually is the following
$\sup_{C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\sum_{c\sim C}\frac{1}{c}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}S(n,\ell;c)V_{c,\ell}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.62)
with
$V_{c,\ell}(x)=V(x)\,e{\left(-\frac{t\log x\mathcal{Y}}{2\pi}+\frac{4\pi\sqrt{x\mathcal{Y}\ell}}{c}\right)}$
for any $x\in\mathbb{R}^{+}$.
Case I: when $C$ is small. Let $C^{\star}\geq 1$ be a parameter determined later. We first consider the case where $C$ is suitably small, such that $C<C^{\star}$. In this case, it is clear that
$\mathcal{T}(K,t)\ll{C^{\star}}^{1/2+\varepsilon}\sqrt{\mathcal{Y}}$ (4.63)
by using the Weil bound for Kloosterman sums.
Case II: when $C$ is large. Now, we turn to the case where $C\geq C^{\star}$. By invoking the Voronoĭ summation formula, Lemma 2.1, this sum is thus essentially turned into
$\begin{split}\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast},r\\ m^{\ast}\geq 1\\ r|c\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{m^{\ast}r^{2}}\,e\left(\mp\frac{m^{\ast}r^{2}\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m^{\ast}r^{2}\mathcal{Y}}{c^{3}};{V}_{c,\ell}}\right).\end{split}$ (4.64)
It suffices to consider the salient scenario where $r=1$, as hinted by the preceding argument in §3.2; the disturbance of the parameter $r$ has no any influence on the final result by an adaptation of the same analysis. The heart of the task thence lies on
$\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast}\geq 1\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{m^{\ast}}e\left(\mp\frac{m^{\ast}\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m^{\ast}\mathcal{Y}}{c^{3}};{V}_{c,\ell}}\right),$
which, by Lemma 2.2, is approximately
$\begin{split}\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\frac{\mathcal{Y}^{1/6}}{C}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast}\geq 1\end{subarray}}\frac{\overline{A_{F}(m^{\ast},1)}}{{m^{\ast}}^{1/3}}\,e\left(\mp\frac{m^{\ast}\overline{\ell}}{c}\right)\,\widehat{V_{\pm,\ell}}\left({\frac{m^{\ast}\mathcal{Y}}{c^{3}}}\right),\end{split}$ (4.65)
with the integral transform $\widehat{V_{\pm,\ell}}$ being
$\widehat{V_{\pm,\ell}}(x)=\int_{\mathbb{R}^{+}}V(y)\,e{\left(-\frac{t\log y}{2\pi}+\frac{4\pi\sqrt{y\mathcal{Y}\ell}}{c}\pm 3(xy)^{1/3}\right)}\mathrm{d}y.$
Lemma 4.1.
Let the integral $\widehat{V_{\pm,\ell}}(x)$ be defined as above. Then we have
  • $\widehat{V_{+,\ell}}(x)\ll\mathcal{Y}^{-A}$ for any sufficiently large $A>0$;
  • $\widehat{V_{-,\ell}}(m^{\ast}\mathcal{Y}/c^{3})$ is negligibly small unless that $m^{\ast}\asymp\sqrt{\mathcal{Y}}\ell^{3/2}$, and in which case it is essentially given by
    $\widehat{V_{-,\ell}}\left({\frac{m^{\ast}\mathcal{Y}}{c^{3}}}\right)\asymp\frac{\sqrt{c}}{(\mathcal{Y}m^{\ast})^{1/6}}\,e{\left(\frac{2\mathcal{Y}^{1/6}{m^{\ast}}^{2/3}}{c\sqrt{\ell}}+O{\left(\frac{(|t|+1){m^{\ast}}^{1/3}}{\mathcal{Y}^{1/6}\sqrt{\ell}}\right)}\right)}.$ (4.66)
Proof.
For any 1/$21/2\leq y\leq 5/$ 2, one defines
$f(y)=\mathcal{A}y^{1/2}\pm\mathcal{B}y^{1/3}-{t}\log y/{2\pi},$ (4.67)
with $\mathcal{A}:=\sqrt{\mathcal{Y}\ell}/c$, and $\mathcal{B}:=3x^{1/3}$. Notice that $\mathcal{A}\geq K^{2}\gg t$. One thus finds that the $+^{\prime}$ case will become negligible, by repeated integration by parts many times. Next, we shall mermly consider the $-^{\prime}$ case, and compute the root of $f^{\prime}(y)=0$; this yields
$\frac{1}{2}\mathcal{A}y^{1/2}-\frac{1}{3}\mathcal{B}y^{1/3}-t/2\pi=0.$ (4.68)
We may approximate the root $y=y^{\flat}$ by writing $y^{\flat}=y_{0}+y_{1}+y_{2}+\cdots$. To be specific, set $y_{0}=\left({2\mathcal{B}}/{3\mathcal{A}}\right)^{6}$ (which follows by equating the first and second terms on the left-hand side of the equation). One computes $y_{1}$ from the equation
$\frac{1}{2}\mathcal{A}\left(y^{1/2}_{0}+\frac{1}{2}y_{1}y^{-1/2}_{0}\right)-\frac{1}{3}\mathcal{B}y^{1/3}_{0}-t/2\pi=0,$
where we have approximated $y^{1/2}$ by $y^{1/2}_{0}+y_{1}y^{-1/2}_{0}/2$, and $y^{1/3}$ by $y^{1/3}_{0}$ in (4.68), respectively. This gives that $y_{1}=\frac{2t}{\pi\mathcal{A}}\sqrt{y_{0}}$. Further, by solving the equation
$\frac{1}{2}{\mathcal{A}}\left(y^{1/2}_{0}+\frac{1}{2}(y_{1}+y_{2})y^{-1/2}_{0}\right)-\frac{1}{3}\mathcal{B}\left(y^{1/3}_{0}+\frac{1}{3}y_{1}y^{-2/3}_{0}\right)-t/2\pi=0,$
we infer that $y_{2}=\frac{\mathcal{B}}{\mathcal{A}}\frac{8t}{9\pi\mathcal{A}}y^{1/3}_{0}$. It follows that $\widetilde{V}_{-,\ell}(x)$, as a function with respect to $x$, is essentially given by the following
$\frac{1}{{\mathcal{B}}^{1/2}}{y}^{-it}_{0}\,e\left(\frac{1}{2}{\mathcal{B}}(y_{0}+y_{1}+y_{2}+\cdots)^{1/3})\right)\approx\frac{1}{{\mathcal{B}}^{1/2}}{y}^{-it}_{0}\,e\left(\frac{2\mathcal{B}^{2}}{9\mathcal{A}}+O{\left(\frac{t{\mathcal{B}}}{\mathcal{A}}\right)}\right).$
Note that we have not taken into account the contribution $e(-t\log y_{0}/2\pi)$ in (4.67), as far as the disturbance is concerned. ∎
As a direct consequence, by Lemma 4.1, one finds that the sum (4.65) is essentially of the fowling form
$\begin{split}&\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\sum_{c\sim C}\frac{1}{\sqrt{c}}\sum_{m^{\ast}\geq 1}\frac{\overline{A_{F}(m^{\ast},1)}}{\sqrt{m^{\ast}}}\,e\left(\frac{m^{\ast}\overline{\ell}}{c}\right)\,\widetilde{{V}}_{c,\ell}\left({\frac{m^{\ast}}{\sqrt{\mathcal{Y}}\ell^{3/2}}}\right),\end{split}$
with $\widetilde{{V}}_{c,\ell}(x)=V(x)e{\left(2\sqrt{\mathcal{Y}\ell}/c\cdot x^{2/3}+O_{c,\ell}(|t|x^{1/3})\right)}$ for any $x\in\mathbb{R}^{+}$. Note that the phase of $\widetilde{{V}}_{c,\ell}$ is roughly $\sqrt{\mathcal{Y}\ell}/c$. One also finds that $\frac{\partial^{i}}{\partial x^{i}}\widetilde{{V}}_{c,\ell}(x)\ll_{i}(\sqrt{\mathcal{Y}\ell}/c)^{i}$ for any $i\geq 0$. The next move is to resort to Lemma 2.3; this would yield the following estimate
$\ll\sup_{C^{\star}\leq C\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\sum_{c\sim C}\frac{\sqrt{\mathcal{Y}\ell}}{c^{3/2}}\cdot\mathcal{Y}^{1/8+\varepsilon}\ell^{3/8}\ll\frac{\mathcal{Y}^{5/8+\varepsilon}\ell^{7/8}}{\sqrt{C^{\star}}},$ (4.69)
where it is necessary that $\ell\gg K^{4}/X$ and $C^{\star}<\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}$ to ensure that the sum over $c$ is not vacuous. Combining with (4.63), it follows that
$\mathcal{T}(K,t)\ll\mathcal{Y}^{9/16+\varepsilon}\ell^{7/16},$
upon choosing $C^{\star}=\mathcal{Y}^{1/8}\ell^{7/8}$. One observes that $C^{\star}<\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}$, provided that $\ell<\mathcal{Y}^{1-\varepsilon}/K^{16/3}$.
4.1.2. Bounding $\widehat{\mathcal{T}}$
Next, in this part let us take a look at the term $\widehat{\mathcal{T}}$ in (4.59). This can be handled by a similar argument. Actually, by repeating the procedure as in §4.1, one sees that $\widehat{\mathcal{T}}$ is boiled down to evaluating
$\begin{split}\sup_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll C\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m\\ m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,1)}}{m}\,e\left(\mp\frac{m\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m\mathcal{Y}}{c^{3}};{U}_{c,\ell}}\right),\end{split}$ (4.70)
with ${U}_{c,\ell}(x)=V(x)e{\left(-{t\log x\mathcal{Y}}/{2\pi}\right)}$ for any $x>0$. By Lemma 2.2, the display above turns out to be controlled by
$\frac{\mathcal{Y}^{1/6}}{C}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,1)}}{{m}^{1/3}}e\left(\mp\frac{m\overline{\ell}}{c}\right)\,\widehat{U_{\pm,\ell}}\left({\frac{m\mathcal{Y}}{c^{3}}}\right),$ (4.71)
where $\widehat{U_{\pm,\ell}}(y)=\int_{\mathbb{R}^{+}}V(x)e(-t\log x/2\pi\pm 3(yx)^{1/3})\mathrm{d}x$ for any $y\in\mathbb{R}^{+}$. Recall that $t\gg\mathcal{Y}^{100\varepsilon}$. The ‘+’ case only survives. In that case, the integral is truncated at $x\asymp|t|^{3}$, and is roughly given by $x^{-1/6}e(\varrho t)V(t^{3}/(8\pi^{2}x))$ for some non-zero constant $\varrho\in\mathbb{R}$. This reveals that the display above is dominated by the upper-bound for
$\begin{split}&\frac{1}{\sqrt{C}}\sum_{c\sim C}\,\sum_{m\asymp C^{3}|t|^{3}/\mathcal{Y}}\frac{\overline{A_{F}(m,1)}}{\sqrt{m}}\,e\left(\frac{m\overline{\ell}}{c}\right)\,V^{\ddagger}\left(\frac{8\pi^{2}m\mathcal{Y}}{C^{3}|t|^{3}}\right),\end{split}$ (4.72)
where $V^{\ddagger}(x)=e{\left(\varrho/(2\pi^{2/3})tx^{1/3}\right)}V(x)$ for any $x\in\mathbb{R}^{+}$, satisfying that $\frac{\partial^{i}}{\partial x^{i}}{V^{\ddagger}}(x)\ll_{j}|t|^{j}$ for any $j\geq 0$. Notice that, here, if $\ell\gg K^{1+\varepsilon}/|t|^{3}$, one has $C^{3}|t|^{3}/\mathcal{Y}\gg 1$. Now, we attempt to apply Lemma 2.3, which shows the estimate that
$\widehat{\mathcal{T}}(K,t)\ll\frac{|t|^{7/4}C^{5/4}}{\mathcal{Y}^{1/4-\varepsilon}}\ll\frac{|t|^{7/4}\mathcal{Y}^{3/8+\varepsilon}\ell^{5/8}}{K^{5/4}}.$ (4.73)
In summary, upon collecting the bounds in (4.60), (4.61), (4.69) and (4.73), we obtain
Proposition 4.2.
For any $\ell\in\mathbb{R}^{+}$ such that
$1+\frac{K^{4+\varepsilon}}{\mathcal{Y}}+\frac{K^{1+\varepsilon}}{|t|^{3}}\leq\ell<\frac{\mathcal{Y}^{1-\varepsilon}}{K^{16/3}},$ (4.74)
with $K$ sufficiently large and $K^{\varepsilon}\ll|t|\ll K^{2-\varepsilon}$, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2+it,F\otimes f)$
$\hskip 65.44142pt\ll\mathcal{Y}^{\varepsilon}\left(\frac{K^{1+\varepsilon}}{\sqrt{\ell}}+\frac{\mathcal{Y}^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}}+\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}}+\mathcal{Y}^{9/16+\varepsilon}\ell^{7/16}+\frac{|t|^{7/4}\mathcal{Y}^{3/8+\varepsilon}{\ell}^{5/8}}{K^{5/4}}\right).$ (4.75)
By inserting the value of $\mathcal{Y}$ and a straightforward simplification, this immediately gives the estimate (1.4), as desired. This assertion of Theorem 1.4 thus follows.
4.2. Proof of Corollary 1.5
At the end of this paper, it remains to complete the proof of Corollary 1.5. Indeed, after divided by a factor $\sqrt{\ell}$ on both sides of (4.2), we sum over $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$ for any $\delta,\varepsilon>0$, obtaining
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2+it,F\otimes g)$
$\hskip 73.97733pt\ll\mathcal{Y}^{\varepsilon}\left(K+\mathcal{Y}^{9/16+\varepsilon}K^{15/16}+\frac{\mathcal{Y}^{3/4}}{K^{7/4}}+\frac{\mathcal{Y}}{K^{3-\varepsilon}}+\frac{|t|^{7/4}\mathcal{Y}^{3/8+\varepsilon}}{K^{1/8}}\right).$
This gives (1.6). The range $K^{10/9+\varepsilon}\leq t\leq K^{2-\varepsilon}$ stems from (4.74), upon taking $\ell$ such that $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$.