### v4

# First moments of ${{\mathrm{GL}}}(3)\times{{\mathrm{GL}}}(2)$ and ${{\mathrm{GL}}}(2)$ $L$-functions and their applications

Abstract.
Let $F$ be a self-dual Hecke-Maaß form for ${\mathrm{GL}}(3)$ underlying the symmetric square lift of a ${\mathrm{GL}}(2)$-newform of square-free level and trivial nebentypus. In this paper, we are interested in the first moments of the central values of $\rm GL(3)\times GL(2)$ $L$-functions and $\rm GL(2)$ $L$-functions. As a result, we obtain an estimate for the first moment for $L(1/2,F\otimes f)$ in a family, where $F$ is of the level $q^{2}$, and $f\in\mathcal{B}^{\ast}_{k}(M)$ for any primes $q,M\geq 2$ such that $(q,M)=1$. We prove the subconvex bound for $L(1/2,F\otimes f)$ involving the levels aspects simultaneously in the range $M^{13/64+\varepsilon}\leq q\leq M^{11/40-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$ for the first time. Moreover, we further investigate the first moments of these $L$-functions in the weight $k$ aspect over $K\leq k\leq 2K$, with $K$ being a large number. As the results, we obtain a Lindelöf average bound for the first moment of $L(1/2,f)L(1/2,F\otimes f)$ of degree 8 and an asymptotic formula for the first moment of $L(1/2,F\otimes f)$ with an error term of $O(K^{-1/4+\varepsilon})$, respectively.
Key words and phrases:
Automorphic forms, First moment, Rankin-Selberg $L$-functions
2010 Mathematics Subject Classification:
11F67, 11F66, 11L05
1. Introduction
The subconvexity problem for $L$-functions is a very basic topic in number theory, which plays an important rôle in arithmetic. In history, for $L$-functions of degree one or two, this problem is completely solved. There are considerable literature on subconvexity of these $L$-functions in various aspects. For degree of three or higher degrees, this becomes very challenging, and only a limited number of results were known.
Typically, a powerful way in studying subconvexity is to appeal to the moment method. See notably Ivić [22], Conrey-Iwaniec [11], Xiaoqing Li [35] and Blomer [2] for relevant descriptions and the references therein. Let $N_{1},N_{2}\geq 1$ be two integers, and $k$ an even integer. Let $F(z)$ be a normalized Hecke-Maaß form of type $\nu=(\nu_{1},\nu_{2})$ for the congruent subgroup $\Gamma_{0}(N_{1})$ with trivial nebentypus, and $f\in\mathcal{B}^{\ast}_{k}(N_{2})$ be a Heke newform on $\text{GL}(2)$ of weight $k$ and level $q$ with trivial nebentypus; see §2.2 for definitions and backgrounds. In this paper, we will be concerned about the central values of $\text{GL}(3)\times\text{GL}(2)$ Rankin-Selberg $L$-function $L(s,F\otimes f)$, and the product of $\text{GL}(3)\times\text{GL}(2)$ and $\text{GL}(2)$ $L$-functions $L(s,F\otimes f)L(s,f)$. This is a very crucial step in understanding automorphic $L$-functions of higher degrees of six and eight.
In 2012, Xiaoqing Li [35] made the first breakthrough to break the convex barriers of $\text{GL}(3)$ $L$-functions in the $t$-aspect as well as $\text{GL}(3)\times\text{GL}(2)$ $L$-functions in the spectral aspect by investigating the first moment of $\text{GL}(3)\times\text{GL}(2)$ $L$-functions. Not long after this, Blomer [2] achieved the subconvexity for $\text{GL}(3)\times\text{GL}(2)\times\text{GL}(1)$ $L$-functions and $\text{GL}(3)\times\text{GL}(1)$ $L$-functions in their conductors aspects. With many sophisticated methods being developed in recent few years such as $\delta$-method (see, for example, [40, 41, 43]), amplification method (see, for example, [19]), deep analysis of the average involving the Kloosterman sums (see, for example, [31, 47]) as well as the integral transforms (see, for example, [41]) to name a few, the researchers were gradually attempting to solve the following well-known subconvexity conjecture:
Conjecture 1.1.
Let $N_{1},N_{2}\geq 1$ be two integers. Let $k$ be an even integer and $t\geq 1$. For any normalized ${{\mathrm{GL}}}(3)$-Maaß form $F$ of level $N_{1}$ and ${{\mathrm{GL}}}(2)$-Maaß form $f$ of level $N_{2}$ and weight $k$, there holds that $L(1/2+it,F\otimes f)\ll_{\varepsilon}(N_{1}N^{3/2}_{2}k^{3}t^{3})^{1/2-\varepsilon}$ for some absolute constant $\varepsilon>0$.
We will collate some results on this conjecture in different regards.
  • In history, Khan [28] firstly broke the convex bound in the $N_{2}$-aspect. In the case of $L(1/2,F\otimes f)\geq 0$, he exploited the amplification method to establish a conditional subconvexity
    $L(1/2,F\otimes f)\ll N^{3/4-1/2001+\varepsilon}_{2},$
    when $N_{2}$ is a prime (which should be though to be the hardest case).
  • Sharma and Sawin [46] employed a type of ‘Mass transfer’ as given by Munshi [42], showing that
    $L(1/2,F\otimes f)\ll p^{3/2-1/16+\varepsilon}$
    for any $f\in\mathcal{B}_{k}(p^{2},\chi^{2})$ with the nebentypus $\chi^{2}$ being of modulus $p^{2}$ (where $p$ is a prime). Later, Kumar, Mallesham and Singh [31] gave a better exponent of $3/2-3/20+\varepsilon$ in the depth aspect, where $f\in\mathcal{B}_{k}(p^{2r},\chi^{2})$ with the nebentypus $\chi^{2}$ being of modulus $p^{r}$, $r>3$.
  • Concerning the $t$-aspect subconvexity for $\text{GL}(3)\times\text{GL}(2)$, Munshi [43] developed the $\delta$-method to its furthest achievement by successfully establishing the following
    $L(1/2+it,F\otimes f)\ll t^{3/2-1/42+\varepsilon}.$
    Subsequently, there were in droves many works devoted to the strengthened (or the generalized) subconvexity; see, for example, [36, 21, 17]. The best exponent should be $3/2-3/20+\varepsilon$ due to Lin and Sun [36].
  • Recently, Kumar [30] firstly proved the subconvex bound in the weight $k$ aspect. By using the ‘conductor lowering mechanism’ developed by Munshi [41], he showed that
    $L(1/2+it,F\otimes f)\ll k^{3/2-1/51+\varepsilon}.$
  • Regarding the level aspect subconvexity involving the $\text{GL}(3)$ forms, it is completly out of reach. By currently developed techniques, it appears that one cannot get any power-saving to beat the subconvex barrier whatsoever. More recently, Kumar, Munshi and Singh [32] were able to prove that
    $L(1/2,F\otimes f)\ll(N_{1}N^{3/2}_{2})^{1/2+\varepsilon}\left(\frac{N^{1/4}_{1}}{N^{3/8}_{2}}+\frac{N^{1/8}_{2}}{N^{1/4}_{1}}\right),$ (1.1)
    when $N_{1},N_{2}$ are both co-prime primes. This hybrid bound is subconvex in the range $N^{1/2+\varepsilon}_{2}<N_{1}<N^{3/2-\varepsilon}_{2}$.
1.1. Main results
In this paper, we are more concerned about the self-dual ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions, which have important implications to equidistribution problems of $L^{q}$-mass of cusp forms (as $q\rightarrow\infty$) and Heegner points on Shimura curves as well as the Quantum Unique Ergodicity (QUE) conjecture (see, for example, [5, 16, 3, 38, 37, 15, 7, 44]). For instance, the well-known Watson’s formula implies that
$\|g\|_{4}^{4}\,\approx\frac{1}{p^{2}}\sum_{h\in\mathcal{B}_{2k}^{\ast}(p)}L(1/2,h)L(1/2,\text{sym}^{2}g\otimes h)$
for any ${\mathrm{GL}}(2)$-newform $g\in\mathcal{B}^{\ast}_{k}(p)$ of large fixed weight $k$ and prime level $p$. Notice that the level aspect subconvexity for self-dual ${\mathrm{GL}}(3)$ forms is still wide-open for specialists, albeit the notable successes of subconvexity for ${\mathrm{GL}}(3)$ $L$-functions in various aspects over the past decade. One of main objectives of the present paper is to address the subconvexity in the levels aspects for ${\mathrm{GL}}(3)$ and ${\mathrm{GL}}(2)$ forms simultaneously [^1] One might wander whether the method of Kumar, Munshi and Singh can be adpated to obtain subconvexity for $L(1/2,\text{sym}^{2}g\otimes f)$ with $F=\text{sym}^{2}g$ being a symmetric square lift of a ${\mathrm{GL}}(2)$-newform $g$, say. Unfortunately, this seems impracticable. The underlying reason is that the modulus $\mathcal{Q}$ in equipping with $\delta$-symbol method is usually large relative to the level of $F$; see [32, Section 2-2B]. This gives rise to a deadlock since of an inevitable gap for ${\mathrm{GL}}(3)$-Voronoĭ formulae which we will explain after a while.. To be specific, we prove the following results:
Theorem 1.2.
Let $q,M\geq 2$ be two primes satisfying that $(q,M)=1$ and $M^{1/8}<q<M^{3/8}$. Let $F$ be a Hecke-Maaß form for the congruence subgroup $\Gamma_{0}(q^{2})$ obtained as the symmetric square lift of a ${\mathrm{GL}}(2)$-newform of level $q$ with trivial nebentypus. Define the set
$\mathscr{P}_{L}=\{p:p\text{ is a prime },(p,qM)=1,L\leq p\leq 2L\}$ (1.2)
with
$\frac{q^{1+\varepsilon}}{{M^{1/4}}}+\frac{{M}^{1/4}}{{q^{1-\varepsilon}}}\leq L\leq M^{1/8-\varepsilon}.$
Then, for any ={1,$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\}$ with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, we have
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f)\ll(qM)^{\varepsilon}\left(\frac{1}{\sqrt{\ell}}+q\sqrt{\frac{\ell}{M}}+\frac{q^{2}\ell}{M^{5/8}}+\frac{q\ell^{2}}{M^{5/8}}+\frac{q^{7/2}\ell^{7/4}}{M^{9/8}}\right),$ (1.3)
where the implied constant depends only on the weight $k$ and the Langlands parameters $\alpha_{i}$, $1\leq i\leq 3$, of the form $F$ as given in (2.4).
Observe that under the self-duality for $\rm GL(3)$ forms, the non-negativity for the central values of the ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions is guaranteed by Lapid’s method [33]; see also [35, Remark 3]. As an application of Theorem 1.2, we are thus able to obtain the following subconvex bound:
Corollary 1.3.
Let the forms $F,f$ and the parameters $q,M,k$ be as in Theorem 1.2. Let $\mathcal{X}\geq 1$ such that $(M^{3/2}q^{2})^{1-\delta}\leq\mathcal{X}\leq(M^{3/2}q^{2})^{1+\varepsilon}$ for arbitrarily small constants $\varepsilon,\delta>0$. Then, for any $L$ with
$\frac{q^{1+\varepsilon}}{{M^{1/4}}}+\frac{M^{(1+3\delta)/4+\varepsilon}}{q^{1-\delta}}\leq L\leq\min{\left\{M^{3/64-\varepsilon},\frac{M^{3/32-\varepsilon}}{q^{1/4}},\,\frac{{M}^{1/8}}{q^{5/14+\varepsilon}},\,\frac{M^{7/40-\varepsilon}}{q^{3/10}}\right\}},$ (1.4)
we have
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2+\varepsilon+\delta}\left(\frac{M^{1/4}}{qL}+\frac{L^{2}}{M^{1/4}}+\frac{L^{4}q}{M^{3/8}}+\frac{L^{8}}{M^{3/8}}+\frac{L^{7}q^{5/2}}{M^{7/8}}+\frac{1}{\sqrt{M}}\right).$ (1.5)
This establishes the subconvexity for any $q,M$ with $M^{13/64+\varepsilon}\leq q\leq M^{11/40-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$, and should be the first instance for the levels aspects subconvexity of self-dual ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions. Taking a little effort, it is not hard to attain the subconvexity bounds for any $q,M$ varying in this range. For instance, when $q\asymp M^{1/4}$, we have
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2-1/80+\varepsilon}$
and when $q\asymp M^{11/40-\varepsilon}$, we obtain
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2-3/82+\varepsilon}.$
In this paper, another motivation is to investigate the first moments of $L(1/2,F\otimes f)$ and $L(1/2,f)L(1/2,F\otimes f)$ in the weight aspect in families. We continue to explore the first moment method by adding a ‘harmonic’ average of the weight $k$, with $K<k\leq 2K$. As a result, we are able to establish the following Lindelöf average bound:
Theorem 1.4.
Let $F(z)$ be a normalized self-dual Hecke-Maaß form on ${{\mathrm{GL}}}(3)$. Let $K$ be a sufficiently large parameter. Let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Then, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2,F\otimes f)\ll K^{1+\varepsilon},$ (1.6)
where the implied constant depends only on the constant $\varepsilon$, the form $F$ as well as the level of the form $f$.
In addition, we are able to obtain the following asymptotic formula for the ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions in the weight aspect with a power-saving error term:
Theorem 1.5.
Keep the notation as in Theorem 1.4. Then, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,F\otimes f)=\frac{L(1,F)K}{4}\widehat{W}(0)+O(K^{-1/4+\varepsilon}),$ (1.7)
where $\widehat{W}$ denotes the ordinary Fourier transform of $W$.
Consequently, as a by-product we can obtain a cute result for the non-vanishing problem of $L(1/2,F\otimes f)$ as follows:
Corollary 1.6.
For any self-dual Maaß from $F$ on ${\mathrm{GL}}(3)$, there exists a newform $f\in\mathcal{B}^{\ast}_{k}$ with $k$ sufficiently large, such that $L(1/2,F\otimes f)\neq 0$.
It is instructive to see that the estimate in (1.6) should be the best-possible in essence, and is already able to beat what the classic spectral large sieve (see, for example, [12, Theorem 2]) together with the Cauchy-Schwarz inequality implies, which only provides a quannity of $O(K^{2+\varepsilon})$. This also reflects the significant effectiveness by further averaging over the parameter $k\sim K$, when it comes to investigating the subconvexity problem involving the weight aspect. We also refer the readers to the works of Luo [39] and Khan [27] for relevant heuristics. It, on the other hand, should be noticed that, thanks to the rational structure of the exponential factor in (4.7), we would succeed to achieve this strong Lindelöf average bound; see Remark 2.10 below for details. Besides, it is also believed that our method in proving Theorem 1.6 can be naturally adapted to produce simultaneous non-vanishing of ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ and ${\mathrm{GL}}(2)$ $L$-functions at the central point, and even subconvexity for $L(1/2,F\otimes f)$ or $L(1/2,f)L(1/2,F\otimes f)$ in the weight and the level (of $F$ or $f$) aspects. We are now working on this direction, and expect to report the progress in the future article.
The initial manoeuvre of our proofs of main results follows the path pioneered by Xiaoqing Li and Blomer, who looked into the average of first moment of $\rm GL(3)\times GL(2)$ $L$-function in a family. The strategies we pursued in the paper, however, can be understood as the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Bilinear form” in proving Theorem 1.2 and “Kuznetsov-Exponential sums/Voronoĭ” in proving Theorems 1.4 & 1.7, respectively. In particular, for the former, this is reminiscent of an important work of Blomer and Khan, [6], in which they investigated the twisted moments of $\text{GL}(3)\times\text{GL}(2)$ and $\text{GL}(2)$ automorphic $L$-functions, and established the following reciprocity formula
$\frac{1}{\sqrt{q}}\sum_{f\text{ level }q}L(1/2,F\otimes f)L(1/2,f)\lambda_{f}(\ell)\,\leftrightsquigarrow\,\frac{1}{\sqrt{\ell}}\sum_{f\text{ level }\ell}L(1/2,F\otimes f)L(1/2,f)\lambda_{f}(q)$ (1.8)
by the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Kuznetsov” law. Among other things, one salient point is the application of the reciprocity law, which, as the Kloosterman sum of dimension one, keeps us going back to somewhere. However, when applying the Voronoĭ to the Kloosterman sum $S(n\ell,1;c)$ of dimension two, the inverse of the variable $\ell$, that is, the exponential factor like $S(n\overline{\ell},\sigma;c)$ for some less significant parameter $\sigma$, will be occurred; this will increase the dimensions of Kloosterman sums afterwards, if one were to resort to the Voronoĭ again. In this sense, we can barely expect the reciprocity formula for Hecke-Maaß forms of higher degrees $d\geq 4$ in the levels aspects [^2] Alternatively, one might explore the reciprocity by equipping with the Kuznetsov formulae in the level aspect for ${\rm{GL}}(n)$, $n\geq 3$; see, for example, [4] for descriptions on ${\rm{GL}}(3)$. However, observe that, it seems very tricky to establish the relation $S(m_{1}m_{2}\cdots m_{k}{c_{1}}\overline{\ell},1;c_{2})\leftrightsquigarrow S(n_{1}n_{2}\cdots n_{k}{c_{1}}\overline{q},1;c_{2})$ for the levels $\ell,q\geq 1$, where $m_{i},n_{i}$ are related to the summations of Fourier coefficients of ${\mathrm{GL}}(d_{i})$-forms, $d_{i}\geq 1$. It is still not sure that if a kind of reciprocity law involving the ${\mathrm{GL}}(3)$ level aspect can be achieved.. It seems that the possible symmetric relation like (1.8) for $\text{GL}(d)\times\text{GL}(2)$ involving the levels aspects is only for the $\text{GL}(3)\times\text{GL}(2)$ case (we refer the interested readers to Jana and Nunes’s recent work [25] for the generalization to $\rm{PGL}_{d+1}\times\rm{PGL}_{d-1}$). In contrast to the $\text{GL}(3)\times\text{GL}(2)$ case, we notice that, for the $\text{GL}(2)\times\text{GL}(1)$ case, Conrey [10] proved the conductors aspects reciprocity formula
$\frac{1}{\sqrt{p}}\sum_{\chi\bmod p}|L(1/2,\chi)|^{2}\chi(\ell)\,\leftrightsquigarrow\,\frac{1}{\sqrt{\ell}}\sum_{\chi\bmod\ell}|L(1/2,\chi)|^{2}\chi(-p).$
This, however, was further extended by Young [49] and Bettin [1] subsequently.
Let us now elucidate two of main obstacles in the present paper. It is remarkable that, by the current development of the Voronoĭ formulae for ${\mathrm{GL}}(3)$ in the level aspect, there is still a gap for arbitrary non-zero integer $c\in\mathbb{Z}$ (see Lemma 2.1 below). If the additive twist colludes with the level $N$ of the form $F$, namely, the parameters $c,N$ satisfy that $(N/(N,c),c)\neq 1$. For instance, one takes $N=q^{2}$ and $c=c^{\prime}q$ with $(c^{\prime},q)=1$ for any prime $q$. The ${{\mathrm{GL}}}(3)$-Voronoĭ formula in this case looks rather tricky. In principle, Corbett’s formula [9, Theorem 1.1] can cover this, but it requires a non-trivial analysis of the $p$-adic Bessel transforms, which, however, becomes more involved; see [9] for relevant details. We also refer the readers to the work of Booker, Milinovich and Ng [8], who worked out the Voronoĭ formula in the ${{\mathrm{GL}}}(2)$-case (if the additive twist colludes with the level). As an innovation in this paper, we are able to overcome this deadlock, which makes the hybrid subconvexity for ${\mathrm{GL}}(3)$ and ${\mathrm{GL}}(2)$ in the levels aspects come into being. The key point, however, is to invoke the reciprocity law as mentioned above. Indeed, with a suitable constraint for the parameter $\ell$ we are in the unramified case, when applying the Voronoĭ for the first time; then, one would be transferred to the ramified case after an implementation of the reciprocity law. This makes it possible for one to equip with the Voronoĭ again.
The second obstacle is the root number. It is well-known that the explicit determination of root numbers play a focal rôle when doing averages of central values of $L$-functions in different families. In this paper, we will give a detailed description of the root number $\varepsilon(F\otimes f)$ for any ${\mathrm{GL}}(3)$ Hecke-Maaß form $F$ with arbitrary level $N$ in the framework of the adelic interpretations, which is of independent interest; see §2.3. Part of descriptions in the special case where $F$ arises as a symmetric square lift from ${\mathrm{GL}}(2)$ has already been given by Holowinsky, Munshi and Qi [18]. As it can be seen later, the root number $\varepsilon(F\otimes f)$ depends only on the parameter $k$ and the form $F$. One of its factor $i^{k}$ will have a direct influence on the ‘harmonic’ averages of the Bessel function $J_{k-1}$ in considering the moments of $L$-values in the weight $k$ aspect; see Lemmas 2.8 and 2.9.
To expose everything as clearly as possible, we have assumed that the parameter $M$ is a prime, which simplifies the argument when applying Petersson trace formula, but our method could in principle be amenable to handling the general case of square-free level with a bit more effort; see the beginning of §3.2 for relevant details.
Acknowledgements. The author is very grateful to Profs. Peter Humphries and Guangshi Lü for helpful discussions during the preparation of this work.
Notations.
Throughout the paper, $\varepsilon$ (resp. $\delta$) always denotes an arbitrarily small positive constant which may not be the same at each occurrence. $A\ll_{\varepsilon}B$ means that $|A|\ll X^{\varepsilon}|B|$, and $n\sim N$ means that $N<n\leq 2N$ for every $n\in\mathbb{Z}$. For any positive integers $m,n$, $(m,n)$ denotes the great common divisor of $m$ and $n$. $\mu$ is the Möbius function.
2. Prerequisites
The purpose of this section is to develop (or supply) certain preliminaries which will be assembled in the bodies of proofs of the main results.
2.1. Automorphic forms
We will give a recap of some fundamental facts about cuspidal forms (see, for example, Iwaniec and Kowalski’s book [23]). Let $k\geq 2$ be an even integer and $M>0$ be a square-free integer. Let $\chi$ be a primitive character of modulus $r$ such that $M|r$, which satisfies that $\chi(-1)=(-1)^{k}$. We denote by $\mathcal{S}_{k}(M,\chi)$ the vector space of holomorphic cuspidal forms on $\Gamma_{0}(M)$ with nebentypus $\chi$ and weight $k$. For any $f\in\mathcal{S}_{k}(M,\chi)$, one then has a Fourier expansion
$f(z)=\sum_{n\geq 1}\psi_{f}(n)n^{\frac{k-1}{2}}e(nz)$
for $z\in\mathbb{H}$. Here, $e(z)$ is shorthand for $e^{2\pi iz}$ for any $z\in\mathbb{C}$, and $\mathbb{H}$ means the upper half-plane. The space $\mathcal{S}_{k}(M,\chi)$ is a finite dimensional Hilbert spaces which can be equipped with the Petersson inner products
$\langle f_{1},f_{2}\rangle=\int_{\Gamma_{0}(N)\backslash\mathbb{H}}f_{1}(z)\overline{f_{2}(z)}y^{k-2}\mathrm{d}x\mathrm{d}y.$
We recall the Hecke operators $\{T_{n}\}$ with $(n,M)=1$, which satisfy the multiplicativity relation
$T_{n}T_{m}=\sum_{d|(n,m)}\chi(d)T_{\frac{nm}{d^{2}}}.$ (2.1)
Moreover, for any $f_{1},f_{2}\in\mathcal{S}_{k}(N,\chi)$, one has
$\langle T_{n}f_{1},f_{2}\rangle=\overline{\chi(n)}\langle f_{1},T_{n}f_{2}\rangle.$
One might thus find an orthogonal basis $\mathcal{B}_{k}(M,\chi)$ of $\mathcal{S}_{k}(M,\chi)$ consisting of common eigenfunctions of all the Hecke operators $T_{n}$ with $(n,N)=1$. For each $f\in\mathcal{B}_{k}(M,\chi)$, denote by $\lambda_{f}(n)$ the $n$-th Hecke eigenvalue, which satisfies
$T_{n}f(z)=\lambda_{f}(n)f(z)$
for all $(n,M)=1$. From (2.1), one has seen that
$\psi_{f}(m)\lambda_{f}(n)=\sum_{d|(n,m)}\chi(d)\psi_{f}\left(\frac{mn}{d^{2}}\right)$
for any $m,n\geq 1$ with $(n,M)=1$. In particular, $\psi_{f}(1)\lambda_{f}(n)=\psi(n)$, if $(n,M)=1$. Therefore, if $(mn,M)=1$,
$\overline{\lambda_{f}(n)}=\overline{\chi(n)}\lambda_{f}(n),\quad\lambda_{f}(m)\lambda_{f}(n)=\sum_{d|(n,m)}\chi(d)\lambda_{f}\left(\frac{mn}{d^{2}}\right).$ (2.2)
The Hecke eigenbasis $\mathcal{B}_{k}(M,\chi)$ also contains a subset of newforms $\mathcal{B}^{\ast}_{k}(M,\chi)$, those forms which are simultaneous eigenfunctions of all the Hecke operators $T_{n}$ for any $n\geq 1$, and normalized to have first Fourier coefficient $\psi_{f}(1)=1$. The elements of $\mathcal{B}^{\ast}_{k}(M,\chi)$ are usually called primitive forms in the sense of Atkin-Lehner; that is, each one is orthogonal to all oldforms, and is an eigenfunction of all the Hecke and Atkin-Lehner operators. For convenience, we abbreviate the notation $\mathcal{B}^{\ast}_{k}(M)$ to $\mathcal{B}^{\ast}_{k}$, the space of newforms of weight $k$ and arbitrary level parameter, if one does not specify the level parameter.
2.2. Automorphic $L$-functions and functional equations
In this part, we shall further shed some light on automorphic $L$-functions. Let $k\geq 2$ be an even integer. We assume that $N\geq 1$ is a positive integer, satisfying that $(N,M)=1$. For any newform $f\in\mathcal{B}_{k}^{\ast}(M)$, we define the attached $L$-function $L(f,s)$, for $\text{Re}(s)>1$, by the Dirichlet series $\sum_{n\geq 1}\lambda_{f}(n)n^{-s}$ which can be expressed in terms of an Euler product
$\prod_{p\nmid M}\left(1-\frac{\lambda_{f}(p)}{p^{s}}+\frac{1}{p^{2s}}\right)^{-1}\prod_{p\mid M}\left(1-\frac{\lambda_{f}(p)}{p^{s}}\right)^{-1}.$
We proceed by defining the completed $L$-function
$\Lambda(s,f)=\mathcal{Q}^{s/2}_{f}L_{\infty}(s,f)L(s,f),$
with
$L_{\infty}(s,f)=\pi^{-s}\Gamma{\left(\frac{s}{2}+\frac{k-1}{4}\right)}\Gamma{\left(\frac{s}{2}+\frac{k-1}{4}\right)}$
and the (analytic) conductor $\mathcal{Q}_{f}$ being $\mathcal{Q}_{f}\asymp M$. The completed $L$-function thus admits an analytical continuation to all $s\in\mathbb{C}$, and satisfies that
$\Lambda(s,f)=\varepsilon(f)\Lambda(1-s,f),$
where the root number $\varepsilon(f)=i^{k}\mu(M)\lambda_{f}(M)\sqrt{M}=\pm i^{k}$.
Let us now put
$\Gamma_{0}(N)=\left\{g\in\text{GL}(3,\mathbb{Z}):g\equiv\left(\begin{matrix}\ast&\ast&\ast\\ \ast&\ast&\ast\\ 0&0&\ast\end{matrix}\right)\mod N\right\}.$ (2.3)
Let $F(z)$ be a normalized Hecke-Maaß form of type $\nu=(\nu_{1},\nu_{2})$ for the congruent subgroup $\Gamma_{0}(N)$ with trivial nebentypus, which has a Fourier-Whittaker expansion with the Fourier coefficients $A_{F}(m,n)$. The Fourier coefficients of $F$ and that of its contragredient $\widetilde{F}$ are related by $A_{F}(m,n)=A_{\widetilde{F}}(n,m)$ for any $(mn,N)=1$, with $A_{F}(1,1)=1$; see, for example, [13, Chapter 6] for definition and backgrounds. The Jacquet-Langlands $L$-function is given by $L(s,F)=\sum_{n\geq 1}A_{F}(n,1)n^{-s}$ for $\text{Re}(s)>1$. Now, let
$\alpha_{1}=-\nu_{1}-2\nu_{2}+1,\quad\alpha_{1}=-\nu_{1}+\nu_{2},\quad\alpha_{1}=2\nu_{1}+\nu_{1}+1$ (2.4)
be the Langlands parameters of $F$. We define the completed $L$-function
$\Lambda(s,F)=\mathcal{Q}^{s/2}_{F}\varepsilon(F)L_{\infty}(s,F)L(s,F),$
where
$L_{\infty}(s,F)=\pi^{-3s/2}\prod_{i=1}^{3}\Gamma{\left(\frac{s+\alpha_{i}}{2}\right)},$
the (analytic) conductor $\mathcal{Q}_{F}\asymp N$, and $|\varepsilon(F)|=1$. One thus have the following functional equation
$\Lambda(s,F)=\varepsilon(F)\Lambda(1-s,\widetilde{F}).$
Next, for any $f\in\mathcal{B}^{\ast}_{k}(M)$, we turn to considering the $\text{GL}(3)\times\text{GL}(2)$ Rankin-Selberg $L$-function $L(s,F\otimes f)$, which is given by the Dirichlet series
$L(s,F\otimes f)=\sum_{m,n\geq 1}\frac{A_{F}(n,m)\lambda_{f}(n)}{(nm^{2})^{s}}$
for $\text{Re}(s)>1$. Now, one defines the Gamma factor $L_{\infty}(s,F\otimes f)$ by
$\begin{split}L_{\infty}(s,F\otimes f)&=\pi^{-3s}\prod_{i=1}^{3}\Gamma{\left(\frac{s+{(k-1)}/{2}+\alpha_{i}}{2}\right)}\Gamma{\left(\frac{s+{(k+1)}/{2}+\alpha_{i}}{2}\right)}\end{split}$ (2.5)
and the (analytic) conductor $\mathcal{Q}_{F,f}\asymp N^{2}M^{3}$. The completed $L$-function defined by
$\Lambda(s,F\otimes f)=\mathcal{Q}^{s/2}_{F,f}L(s,F\otimes f)$
is thus an entire function with an analytic continuation to all $s\in\mathbb{C}$, and satisfies the function equation that
$\Lambda(s,F\otimes f)=\varepsilon(F\otimes f)\Lambda(1-s,\widetilde{F}\otimes f).$
2.3. Epsilon constant
In this part, we will be dedicated to developing some facts on the epsilon constant $\varepsilon(F\otimes f)$, which play a key rôle in later analysis. We proceed by borrowing some points of view from the automorphic representations of algebraic groups over local fields. Let the parameters $k,N,M$ be as before. Define
$k_{2,p}=\left\{g\in\text{GL}(2,\mathbb{Z}_{p}):g\equiv\left(\begin{matrix}\ast&\ast\\ 0&\ast\end{matrix}\right)\bmod p\right\}$
and
$k_{3,p}=\left\{g\in\text{GL}(3,\mathbb{Z}_{p}):g\equiv\left(\begin{matrix}\ast&\ast&\ast\\ \ast&\ast&\ast\\ 0&0&\ast\end{matrix}\right)\bmod p\right\}.$
For any given holomorphic newform $f\in L^{2}(\Gamma_{0}(M)\backslash\text{SL}(2,\mathbb{R}))$ of weight $k$, let us proceed by defining a function $\varphi_{f}\in L^{2}(\text{GL}(2,\mathbb{Q})\backslash\text{GL}(2,\mathbb{A}_{\mathbb{Q}})/\text{K}(2,M))$ by
$\varphi_{f}(\gamma g_{\infty}k)=f(g_{\infty})$
for $\gamma\in\text{GL}(2,\mathbb{Q})$, $g_{\infty}\in\text{GL}(2,\mathbb{R})$ and $k\in\text{K}(2,M)$, with
$\text{K}(2,M)=\text{O}(2,\mathbb{R})\times\prod_{p|M}k_{2,p}\times\prod_{p\nmid M}\text{GL}(2,\mathbb{Z}_{p}).$
It can be seen that each function $\varphi_{f}$ generates the automorphic representation $\pi_{f}$ of $\text{GL}(2,\mathbb{A}_{\mathbb{Q}})$ with trivial central character. We refer the reader to, for example, [14, Sections 13.3-13.6] for relevant backgrounds.
For any given Hecke-Maaß form $F$ of level $N$ on $\text{GL}(3)$, we now let $\Pi_{F}$ be the associated automorphic representation, which belongs to the space $L^{2}(\Gamma_{0}(N)\backslash{\text{SL}}(3,\mathbb{R}))$. Here, the definition of $\Gamma_{0}(N)$ can be referred to (2.3) in §2.1. The representation $\Pi_{F}$ is generated by the smooth functions $\varphi_{\Pi_{F}}\in L^{2}(\textrm{GL}(3,\mathbb{Q})\backslash\textrm{GL}(3,\mathbb{A}_{\mathbb{Q}})/\text{K}(3,N))$, which are defined in the following way
$\varphi_{\Pi_{F}}(\gamma g_{\infty}k)=F(g_{\infty})$
for any $\gamma\in\text{GL}(3,\mathbb{Q})$, $g_{\infty}\in\text{GL}(3,\mathbb{R})$ and $k\in\text{K}(3,N)$, with
$\text{K}(3,N)=\text{O}(3,\mathbb{R})\times\prod_{p|N}k_{3,p}\times\prod_{p\nmid N}\text{GL}(3,\mathbb{Z}_{p}).$
It is well-known that one might factor
$\Pi_{F}=\Pi_{F,\infty}\otimes_{p<\infty}\Pi_{F,p},$
where $\Pi_{F,p}$ is an irreducible, admissible and unitary representations of $\text{GL}(3,\mathbb{Q}_{p})$ for every $p<\infty$.
We shall now ready to deduce an explicit form of the epsilon constant $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})$ at each place $p\leq\infty$. Observe that, for any unramified places $v=p$ with $p\nmid NM$, there already hold that $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=1$. We will focus on the archimedean place $v=\infty$ and the finite places $v=p,p|NM$ in the following.
  • To begin with, at archimedean place $p=\infty$, in the Hecke-Maaß form setting $\Pi_{F,\infty}$ is an unramified principal series representation; while, $\pi_{f,\infty}$ is a ramified discrete series representation of weight $k$, by the multiplicativity of epsilon factors we have
    $\varepsilon(\Pi_{F,\infty}\otimes\pi_{f,\infty})=\varepsilon(\pi_{f,\infty})^{3}=i^{k}.$
  • Now, for $p|M$, the local component $\Pi_{F,p}$ is an unramified principal series representation, and $\pi_{g,p}$ is the Steinberg representation, one finds
    $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=\varepsilon(\pi_{f,p})^{3}=-\lambda_{f}(p)\sqrt{p}.$
  • For $p|N$, one observes that the local component $\Pi_{F,p}$ is the Steinberg representation; whilst, $\pi_{g,p}$ is an unramified principal series representation with trivial central character, so that
    $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=\varepsilon(\Pi_{F,p})^{2},$
    with $|\varepsilon(\Pi_{F,p})|=1$.
We finally obtain
$\varepsilon(F\otimes f)=\prod_{p\leq\infty}\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=-i^{k}\lambda_{f}(M)\sqrt{M}\prod_{p|N}\varepsilon(\Pi_{F,p})^{2}.$ (2.6)
2.4. Voronoĭ summation formula
Let $w$ be a compactly supported smooth function on $(0,\infty)$, and $\widetilde{w}$ be its Mellin transform. For any $\rho=0,1$ and any Hecke-Maaß form $F$ of level $N$ with the Langlands parameters $\alpha_{i}$, $1\leq i\leq 3$, being as in (2.4), we define
$\gamma_{\rho}(s)=\varepsilon(F)i^{\rho}\pi^{3(s-1/2)}\prod_{j=1}^{3}\frac{\Gamma\left(\frac{1+s+\alpha_{j}+\rho}{2}\right)}{\Gamma\left(\frac{-s-\alpha_{j}+\rho}{2}\right)}$ (2.7)
and set
$\gamma_{\pm}(s)=\frac{1}{2}\left(\gamma_{0}(s)\mp i\gamma_{1}(s)\right).$
Now, we proceed to define two auxiliary quantities
$\Omega_{\pm}(x;w)=\frac{1}{2\pi i}\int_{(\sigma)}x^{-s}\gamma_{\pm}(s)\widetilde{w}(-s)\mathrm{d}s,$ (2.8)
where $\sigma>-1+\max\{-\text{Re}(\alpha_{1}),-\text{Re}(\alpha_{2}),-\text{Re}(\alpha_{3})\}$. We then have the following Voronoĭ summation formula for Hecke-Maaß forms on $\text{GL}(3)$ in the level aspect (see [50]).
Lemma 2.1.
Let $w(x)\in\mathcal{C}^{\infty}_{c}(0,\infty)$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Let $a,\overline{a},c\in\mathbb{Z}$ with $c\neq 0$, $(a,c)=1$, satisfying that $a\overline{a}\equiv 1\bmod c$.
  • If $(cm,N)=1$, we then have
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{n\overline{a}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sqrt{N}\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am\overline{N},\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}Nm};w\right).\end{split}$ (2.9)
  • If $N|c$, we, however, have
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{n\overline{a}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am,\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}m};w\right).\end{split}$ (2.10)
One has a good control of the asymptotic behavior of $\Omega_{\pm}$, upon expressing it in term of the oscillatory integral; see, for example, [34, Lemma 6.1].
Lemma 2.2.
For any fixed integer $\mathcal{K}\geq 1$ and $x\gg 1$, we have
$\begin{split}&\Omega_{+}(x;w)=x\int_{\mathbb{R}^{+}}w(y)\sum_{j=1}^{\mathcal{K}}\frac{1}{(xy)^{{j}/{3}}}\left[c_{j}\,e{\left(3(xy)^{1/{3}}\right)}\right.\\ &\phantom{=\;\;}\left.\hskip 163.03416pt+d_{j}\,e{\left(-3(xy)^{1/{3}}\right)}\right]\mathrm{d}y+O{\left(x^{({-\mathcal{K}+2})/{3}}\right)},\end{split}$
where $c_{j},d_{j}$ are suitable constants depending on the four parameters $\alpha_{i}$, $1\leq i\leq 3$, above. Furthermore, $\Omega_{-}(x;\omega)$ has the same expression except values of constants.
This, in turn, reduces the right-hand side of the Voronoĭ formula (2.9) essentially to
$\frac{X^{{2}/{3}}}{cN^{1/6}m^{2/3}}\sum_{\pm,n_{1}|c}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{(n_{2}/n_{1})^{1/{3}}}\,S(am\overline{N},\pm n_{2};cm/n_{1})$
$\hskip 179.25244pt\int_{\mathbb{R}^{+}}w(y)\,e\left(\pm\frac{3\left(n_{2}n^{2}_{1}Xy/m\right)^{1/{3}}}{c}\right)\mathrm{d}y.$ (2.11)
The main contribution occurs when $n_{2}\ll c^{3}m/n^{2}_{1}X^{1-\varepsilon}$, by repeated integration by parts, and in which case $\Omega_{\pm}\ll X^{\varepsilon}$.
2.5. Non-linear exponential sums
We record the following square-root cancellation estimate for the non-linear exponential sums related to Hecke-Maaß forms on $\text{GL}(3)$; see [45, Theorem 3] and the remarks after [45, Eqn. (3.18)].
Lemma 2.3.
Let $X\geq 2$. Let $W$ be a smooth function supported on $[1/2,5/2]$ with partial derivatives satisfying $x^{i}W(x/X)\ll_{i}ZP^{i}$ for any $i\geq 0$ and some $Z,P>0$. Then, for any $\alpha\in\mathbb{R}$ with $\alpha^{2}\in\mathbb{Q}$, we have
$\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}e(\alpha\sqrt{m})\,W{\left(\frac{m}{X}\right)}\ll ZP(1+|\alpha|\log X).$
2.6. Petersson trace formula
For any integers $m,n\geq 1$, we set
$\Delta_{k,M}(m,n)=\sum_{f\in\mathcal{B}_{k}(M)}\omega^{-1}_{f}\psi_{f}(m)\psi_{f}(n)$
and
$\Delta^{\ast}_{k,M}(m,n)=\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(m)\lambda_{f}(n),$
where the spectral weights $\omega_{f}$ are given by $\omega_{f}:=\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\langle f,f\rangle$. In particular, if $f\in\mathcal{B}^{\ast}_{k}(M)$, one has $\omega_{f}=\frac{(k-1)M}{2\pi^{2}}L(1,\text{sym}^{2}f),$ which satisfies that $k^{1-\varepsilon}M\ll\omega_{f}\ll k^{1+\varepsilon}M$ for any $\varepsilon>0$; see [24, Lemma 2.5].
Lemma 2.4.
(Petersson trace formula) We have
$\Delta_{k,M}(m,n)=\delta(m,n)+2\pi i^{-k}\sum_{c>0}\frac{S(m,n;cM)}{cM}J_{k-1}\left(\frac{4\pi\sqrt{mn}}{cM}\right),$
where $\delta(m,n)=1$ if $m=n$, and $\delta(m,n)=0$ otherwise.
For the purpose of our application, we need an expression for $\Delta^{\ast}_{k,M}$ instead of $\Delta_{k,M}$. Suppose that $m,n$ are positive integers satisfying $(m,M)=1$, we have the following identity (see [24, Eqn. (2.51)]):
Lemma 2.5.
We have
$\Delta^{\ast}_{k,M}(m,n)=\sum_{RS=M}\frac{\mu(S)}{L\nu((m,S))}\sum_{l|S^{\infty}}l^{-1}\sum_{l^{2}_{1}|(n,\,Sl_{1})}\mu(l_{1})l_{1}\Delta_{k,R}\left(ml^{2},\frac{n}{l^{2}_{1}}\right),$
where $\nu(M)=[\Gamma(1):\Gamma(M)]=M\prod_{p|M}(1+p^{-1})$.
2.7. Bounds on bilinear forms with Kloosterman sums
The bilinear form with Kloosterman sums always plays a vital rôle in many topics such as the moments of central values of $L$-functions, subconvexity even the classical bilinear exponential sums in additive number theory; see, for example, [31, 26, 47] for descriptions in different regards. In the present paper, we will develop the following relevant results to facilitate our use after a while.
Lemma 2.6.
Let $X,Y\geq 1$. Let $W$ be a smooth function supported on $[1/2,5/2]\times[1/2,5/2]$ with partial derivatives satisfying $x^{i}y^{j}\frac{\partial^{i}}{\partial x^{i}}\frac{\partial^{j}}{\partial y^{j}}W(x/X,y/Y)\ll_{i,j}ZZ^{i}_{1}Z^{j}_{2}$ for some $Z,Z_{1},Z_{2}>0$. Let $q\geq 1$ be an integer. For any $h\in\mathbb{Z}$, we have
$\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(nh,m;q)\,W{\left(\frac{n}{X},\frac{m}{Y},\right)}\ll Z_{1}Z\sqrt{XYq}+Z_{1}ZXq^{3/4}(h,q)^{1/4}+Z_{1}Zq.$ (2.12)
Proof.
This lemma essentially follows from the work of Kerr, Shparlinski, Wu and Xi [26] which appealed to the approach in [47, Theorem 2.1]. By partial summation, one finds that
$\begin{split}&\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(n,hm;q)W{\left(\frac{n}{X},\frac{m}{Y}\right)}\\ &\hskip 85.35826pt\ll\int^{5X/2}_{X/2}\left|\hskip 5.0pt\sideset{}{{}^{\ast}}{\sum}\limits_{x\bmod q}\sum_{m\geq 1}\alpha_{m,t}\,e{\left(\frac{m\overline{x}}{q}\right)}\sum_{n\leq t}e{\left(\frac{nhx}{q}\right)}\right|\mathrm{d}t,\end{split}$
with $\alpha_{m,t}=\frac{\partial}{\partial t}W(t/X,m/Y)A_{F}(m,1)/\sqrt{m}$. This is compatible with the setting as presented in the beginning of the [47, Section 5]. We still might bound the inner triple sum by ±$\sum_{\pm}\sum_{1\leq i\leq I}|S^{\pm}_{i}|$, where, for $1\leq i\leq I$, $S^{\pm}_{i}=\sum_{m\leq X}\alpha_{m,t}\gamma_{x}\,e{({\pm m\overline{x}}/{q})}$; see [26, Section 6] for details. ∎
Observe that the bilinear forms we are concerned about is a smooth-weighted sum. A direct use of the Possion might show some bonus. Indeed, we alternatively have
Lemma 2.7.
Keep the notation as in Lemma 2.6. Then, for any $(h,q)=1$, we have
$\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(nh,m;q)W{\left(\frac{n}{X},\frac{m}{Y}\right)}\ll Z\sqrt{XYq}+\sqrt{Z_{2}}ZXq^{3/4}\mathbf{1}_{Z_{2}q>Y}+ZX\sqrt{Y}.$ (2.13)
Proof.
By switching the orders of the summations and applying the Cauchy-Schwarz inequality followed by the Possion, one deduces that the double sum is controlled by
$\begin{split}&\ll\left(\frac{Y}{q}\sum_{n_{1},n_{2}\geq 1}\,\sum_{|\jmath|\ll Z_{2}q/Y^{1-\varepsilon}}\Phi(\jmath;n_{1},n_{2})\sum_{\alpha\bmod q}S(n_{1}h,\alpha;q)\overline{S(n_{2}h,\alpha;q)}\,e{\left(\frac{\jmath\alpha}{q}\right)}\right)^{1/2},\end{split}$
with $\Phi(\jmath;v_{1},v_{2})=\int_{\mathbb{R}^{+}}W(v_{1},t)W(v_{2},t)e(-\jmath t/q)\mathrm{d}t$ for any $v_{1},v_{2}\in\mathbb{R}^{+}$ and $\jmath\in\mathbb{R}$. It is verifiable that the zero-frequence shows a quantity $Z\sqrt{XYq}+XZ\sqrt{Y}$; while, if $Z_{2}q>Y$, the non-zero frequences, however, will contribute the quantity $\sqrt{Z_{2}}Zq^{3/4}\sqrt{Y}$ by invoking the estimate for triple sums of Kloosterman sums; see, for example, [26, Lemma 4.1]. ∎
It is remarkable that, as it can be seen in (2.13), the upper-bound becomes smaller, as $m,n$ vary in the ranges with relatively small heights.
2.8. Bessel functions
For any $x\geq 1$ and $k\in\mathbb{N}^{+}$, usually one has the crude bound for the Bessel function $J_{k}(x)$ that
$J_{k}(x)\ll{\min{\left(x^{k},x^{-1/{2}}\right)}};$ (2.14)
see, for example, [15, Proposition 8]. For any $x\in\mathbb{R}$, one might also write
$J_{k-1}(x)=e(x)\mathscr{W}^{+}_{x}(s)+e(-x)\mathscr{W}^{-}_{k}(x),$ (2.15)
with
$\mathscr{W}^{\pm}_{k}(x)=\frac{e^{\mp\left(k/2+1/{4}\right)\pi i}}{\sqrt{2\pi x}\,\Gamma\left(k+1/{2}\right)}\int_{\mathbb{R}^{+}}e^{-t}\left(x\left(1\pm\frac{it}{2x}\right)\right)^{k-1/{2}}\mathrm{d}t,$
which satisfy that
$x^{j}{\mathscr{W}^{\pm}_{k}}^{(j)}(x)\ll_{j,k}\min\left(\frac{1}{\sqrt{x}},x^{k-1}\right)$ (2.16)
for any $j\geq 0$; see, for example, [48, Sections 16.12, 17.5 & 16.3]. Now, for any $h$, a smooth bump function supported on $[1/2,5/2]$ with bounded derivatives, we define
[…]
We thus have (see [24, Corollary 8.2]):
Lemma 2.8.
For any $K>0$ and $x>0$, there holds that
[…]
where the implied constant depends only on $h$.
[…]
Lemma 2.9.
For $\iota=0,2$ and any $K>0$ and $x>0$, there holds that
[…]
where the implied constants depend only on $h$.
Remark 2.10.
Observe that the denominator $2\pi$ in both exponential factors of (2.18) and (2.19) usually looks inconspicuous in practical use. However, it plays a miraculous rôle in this paper, and would make it possible for us to apply the square-root estimate in Lemma 2.3. Otherwise, one cannot obtain the Lindelöf average bound in (1.6) whatsoever; see (4.7) and (4.11) below for details.
3. Proofs of Theorem 1.2 and Corollary 1.3
In this section, we are ready to deduce Theorem 1.2, by which one can further show Corollary 1.3.
3.1. Proof of Theorem 1.2
Let $q$ be a prime and $M$ is also, satisfying that $(q,M)=1$. For any $f\in\mathcal{B}^{\ast}_{k}(M)$ and $F$ being a normalized $\text{GL}(3)$ Hecke-Maaß form of level $q^{2}$ with trivial nebentypus, we will consider the first moment of the central value $L(1/2,F\otimes f)$ in the levels aspects. Recall that $\mathcal{Q}_{F,f}\asymp M^{3}q^{4}$. The standard functional equation method (see, for example, [23, Theorem 5.3]) implies that the central value $L(1/2,F\otimes f)$ hinges on the following two quantities
$\begin{split}&\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}U{\left(\frac{n}{q^{2}M^{3/2}}\right)}\quad\text{and}\quad\varepsilon(F\otimes f)\sum_{n\geq 1}\frac{A_{\widetilde{F}}(n,1)\lambda_{f}(n)}{\sqrt{n}}V{\left(\frac{n}{q^{2}M^{3/2}}\right)},\end{split}$
up to two smooth weight functions $U,V$ satisfying $U^{{j}}(x),{V}^{{j}}(x)\ll(1+|x|)^{-A}$ for any $j>0$ and sufficiently large $A>0$. Let $L\geq 1$ be a parameter to be optimized later, satisfying that
$L<M^{1/8}$ (3.1)
and
$L\geq\frac{q^{1+\varepsilon}}{{M^{1/4}}},\quad L\geq\frac{{M}^{1/4}}{{q^{1-\varepsilon}}}$ (3.2)
for any $\varepsilon>0$. We now define a set of modulo
$\mathscr{P}_{L}=\{p:p\text{ is prime, }L\leq p\leq 2L,(p,qM)=1\}.$ (3.3)
Let $f_{0}\in\mathcal{B}^{\ast}_{k}(M)$. By equipping with the amplifier (as in [6, Section 12.2])
$\mathfrak{A}_{f}=\sum_{j=1}^{2}\left|\sum_{\ell\in\mathscr{P}_{L}}\lambda_{f}(\ell^{j})x(\ell^{j})\right|^{2},$ (3.4)
with $x(n)={\text{sgn}}(\lambda_{f_{0}}(n))$, one attempts to estimate an amplified first moment
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\mathfrak{A}_{f}L(1/2,F\otimes f).$
It is clear that, by the relation $\lambda^{2}_{f_{0}}(p)-\lambda_{f_{0}}(p^{2})=1$, one sees that
$\mathfrak{A}_{f_{0}}\gg\frac{1}{2}\left(\sum_{\ell\in\mathscr{P}_{L}}|\lambda_{f_{0}}(\ell)|+|\lambda_{f_{0}}(\ell^{2})|^{2}\right)^{2}\gg\frac{L^{2}}{\log L}.$ (3.5)
If one opens the square and applies multiplicativity of Hecke eigenvalues, finally our main focus is on
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f),$ (3.6)
in view of the fact that
$\mathfrak{A}_{f}=\sum_{\ell\in\mathscr{P}_{L}}\left(x^{2}(\ell)+x^{2}(\ell^{2})\right)$ $+\sum_{\ell_{1},\ell_{2}\in\mathscr{P}_{L}}x(\ell^{2}_{1})x(\ell^{2}_{2})\lambda_{f}(\ell^{2}_{1}\ell^{2}_{2})$
$+\sum_{\ell_{1},\ell_{2}\in\mathscr{P}_{L}}\left(x(\ell_{1})x(\ell_{2})+x(\ell^{2}_{1})x(\ell^{2}_{2})\mathbf{1}_{\ell_{1}=\ell_{2}}\right)\lambda_{f}(\ell_{1}\ell_{2}).$ (3.7)
Notice that, here and thereafter,
={1,$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\},$ (3.8)
with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, is either 1 or a square-free integer or a square satisfying the coprimalities with the parameters $q$ and $M$, upon recalling (3.1).
To proceed, let us recall that the root number $\varepsilon(F\otimes f)=\pm\prod_{p|q}\varepsilon(\Pi_{F,p})^{2}$ according to the foregoing descriptions in §2.3. In the $+^{\prime}$ case, one finds that, the quantity in (3.6) is boiled down to the analysis of the following sum
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}U{\left(\frac{n}{\mathcal{X}}\right)},$ (3.9)
with $\mathcal{Q}^{(1-\delta)/2}_{F,f}\leq\mathcal{X}\leq\mathcal{Q}^{(1+\varepsilon)/2}_{F,f}$ for any small constants $\delta,\varepsilon>0$. Here, $U$ is a smooth bump function supported $[1/2,5/2]$, identical to one for $x\in[1,2]$, with derivatives satisfying that $x^{j}U^{(j)}\left({x}/{\mathcal{X}}\right)\ll_{j}1$ for any $j\geq 0$.
By the Cauchy-Schwarz inequality, together with the Rankin-Selberg bounds for the second moments of Fourier coefficients, the trivial estimate for the double sum in (3.9) is $\ll_{\varepsilon}\mathcal{X}^{1/2+\varepsilon}$. This is far away from the bound $\ll\mathcal{X}^{1/2+\varepsilon}/M$ leading to the convexity. The task is thus reduced to estimating (3.9) further. Our strategy next is the Petersson trace formula. To set up for our application of this, we trivially write the expression (3.9) above as
$\sum_{\begin{subarray}{c}n\geq 1\\ (n,M)=1\end{subarray}}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta^{\ast}_{k,M}(n,\ell)U{\left(\frac{n}{\mathcal{X}}\right)}+\sum_{\begin{subarray}{c}n\geq 1\\ M|n\end{subarray}}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta^{\ast}_{k,M}(n,\ell)U{\left(\frac{n}{\mathcal{X}}\right)}.$ (3.10)
It is straightforward to verify that, by the fact $|\lambda_{f}(M)|=M^{-1/2}$, the second term is bounded by $\ll_{\varepsilon}\sqrt{\mathcal{X}}/M^{3/2}$. For $(n\ell,M)=1$, notice that, by Lemma 2.6 one has the following relation
$\Delta^{\ast}_{k,M}(n,\ell)=\Delta_{k,M}(n,\ell)-\frac{1}{M}\sum_{\rho|M^{\infty}}\frac{1}{\rho}\Delta_{k,1}{\left(n,\ell\rho^{2}\right)},$
so that one may plug the second term on the right-hand side above into (3.10) showing that
$\frac{1}{M}\sum_{\begin{subarray}{c}\rho|M^{\infty}\\ \rho>1\end{subarray}}\frac{1}{\rho}\sum_{\begin{subarray}{c}n\geq 1\\ (n,M)=1\end{subarray}}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}\Delta_{k,1}{\left(n,\ell\rho^{2}\right)}U{\left(\frac{n}{\mathcal{X}}\right)}\ll_{\varepsilon}\frac{\sqrt{\mathcal{X}}}{M^{2-\varepsilon}}.$
This reduces (3.9) to
$\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\left\{\Delta_{k,M}(n,\ell)-\frac{1}{M}\Delta_{k,1}(n,\ell)\right\}U{\left(\frac{n}{X}\right)}$ (3.11)
plus an acceptable error $O_{\varepsilon}{(\mathcal{X}^{1/2+\varepsilon}/M^{3/2})}$. Here, the first term in the braces should be viewed as an average over all forms of level $M$. The second, as an average over the old forms, however, will not contribute significantly compared with the newforms. Indeed, in this case one finds that
$\begin{split}&\frac{1}{M}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta_{k,q}(n,1)U{\left(\frac{n}{X}\right)}=\frac{1}{M}\sum_{f^{\star}\in\mathcal{B}^{\ast}_{k}(1)}\omega^{-1}_{f^{\star}}\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f^{\star}}(n)}{\sqrt{n}}U{\left(\frac{n}{X}\right)},\end{split}$
which is $\ll q^{1+\varepsilon}/M$ by the convexity bound $O(q^{1+\varepsilon})$ for the Rankin-Selberg $L$-function $L(s,F\otimes f^{\star})$. Our job thus reduces to
$\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\Delta_{k,M}(n,1)+O{\left(\frac{\sqrt{\mathcal{X}}}{M^{3/2-\varepsilon}}+\frac{q^{1+\varepsilon}}{M}\right)}.$ (3.12)
Now, we are ready to apply the Petersson trace formula (Lemma 2.4) to $\Delta_{k,M}$; this produces the diagonal and off-diagonal terms of the form $\mathcal{O}(M,q)+2\pi i^{k}\mathcal{S}(M,q)$, where
$\mathcal{O}(M,q)=\sum_{n=1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}\delta(n,\ell)U{\left(\frac{n}{\mathcal{X}}\right)}$ (3.13)
which is $O(\ell^{-1/2+\varepsilon})$, and
$\mathcal{S}(M,q)=\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}U{\left(\frac{n}{\mathcal{X}}\right)}\sum_{c\geq 1}\frac{S(n,\ell;cM)}{cM}J_{k-1}\left(\frac{4\pi\sqrt{n\ell}}{cM}\right).$ (3.14)
Noticing that we are taking $k$ sufficiently large throughout the paper, one finds that the sum over $c$ can be truncated at $c\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$; the contribution in the complementary range is negligibly small (i.e., $O_{A}(\mathcal{X}^{-A})$ for any sufficiently large $A$), with the help of the bound (2.14). It is instructive to see that this sum of $c$ is not vacuous by the second pre-condition in (3.2); otherwise, one cannot succeed in obtaining hybrid subconvexity). In particular, whenever $\ell=1$, it requires that $q\geq M^{1/4+\varepsilon}$ (so that $\sqrt{\mathcal{X}}\gg M^{1+\varepsilon}$ for any $\varepsilon>0$), otherwise the display (3.14) is negligibly small; we will henceforth assume this pre-condition once we are considering the case of $\ell=1$ in the following paragraphs.
By using the dyadic subdivision, we further attempt to estimate the multiple sum for (3.14) as follows
$\sup_{C\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c\sim C}\frac{S(n,\ell;cM)}{cM}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.15)
where ${U}_{\ell}\left(y\right)=U(y)J_{k-1}\left({4\pi\sqrt{y\mathcal{X}\ell}}/{CM}\right)$ for any $y\in\mathbb{R}^{+}$. According to the properties of the $J$-Bessel functions in §2.1, it follows that
$y^{i}\ell^{j}\frac{\partial^{i}}{\partial y^{i}}\frac{\partial^{j}}{\partial\ell^{j}}U_{\ell}(y)\ll\left(\frac{\sqrt{\mathcal{X}L}}{CM}\right)^{i-1/2}$ (3.16)
for any $i,j\geq 0$.
Next, we prepare to apply the Voronoĭ summation formula, Lemma 2.1. Before that, one needs to detect the coprimality between $\ell$ and the modulus of the twist $cM$. Recall that ={1,$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\}$ given as in (3.8).
Case I: $\ell=\ell_{1}\ell_{2}$. To begin with, let us consider what happens whenever $\ell=\ell_{1}\ell_{2}$. It is verifiable that the only counterexample with $(c,\ell)\neq 1$ occurs, provided that $q>M^{1/4+\varepsilon}$, and in which case, $c=\ell_{1}c^{\prime}$ or $\ell_{2}c^{\prime}$ for some $c^{\prime}\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}$, with $(c^{\prime},\ell_{1}\ell_{2}M)=1$ by first one of the restrictions in (3.2). If so, one finds that $S(n,\ell_{1}\ell_{2};\ell_{1}c^{\prime}M)=-S(n,\ell_{2}\overline{\ell_{1}};c^{\prime}M)$, provided that $(n,\ell_{1})=1$, and $S(n,\ell_{1}\ell_{2};\ell_{1}c^{\prime}M)=(\ell_{1}-1)S(n/\ell_{1},\ell_{2};c^{\prime}M)$ otherwise, where $\ell_{1}|n$. Our object (3.15) is essentially degenerated into two terms
$\frac{1}{L}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\prime}\sim C}\frac{S(n,\ell_{2}\overline{\ell_{1}};c^{\prime}M)}{c^{\prime}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.17)
$\frac{A_{F}(\ell_{1},1)}{\sqrt{L}}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\prime}\sim C}\frac{S(n,\ell_{2};c^{\prime}M)}{c^{\prime}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}/L}\right);$ (3.18)
whilst, the expression of (3.15) we are concerned about is exactly
$\sup_{C\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c\sim C}\frac{S(n,\ell_{1}\ell_{2};cM)}{cM}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right).$ (3.19)
It suffices to evaluate (3.19) with $(c,\ell_{1}\ell_{2})=1$, observing that the expression (3.19) bears some resemblances to that in (3.17) and (3.18). By a similar argument, the two possibly-existing quantities (3.17) and (3.18) will provide a relative less significant contribution. For the sake of having a more clear understanding of the two quantities above for readers, we re-affirm that these two terms will not come into being, whenever $q<M^{1/4}$.
Case II: $\ell=\ell^{2}_{1}\ell^{2}_{2}$. Now, we turn to examining that the case of $\ell$ being a square in (3.15). Observe that, the (possibly occurred) scenarios where $(c,\ell)\neq 1$ are as follows:
  • (a)
    $\ell=\ell^{2}_{1}\ell^{2}_{2},c=c^{\ast}\ell_{1}$ or $c^{\ast}\ell_{2}$ for some $c^{\ast}\geq 1$, such that $(c^{\ast},\ell_{1}\ell_{2}M)=1$ and $c^{\ast}\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}$;
  • (b)
    $\ell=\ell^{2}_{1}\ell^{2}_{2},q>M^{1/4+\varepsilon}$, and $c=\hat{c}\ell_{1}\ell_{2}$ for some $\hat{c}\geq 1$, satisfying that $(\hat{c},\ell_{1}\ell_{2}M)=1$ and $\hat{c}\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}$.
Notice that, here, $\ell^{2}_{1}\ell_{1}^{2}|c$ and $\ell^{2}\ell^{2}_{2}|c$ can not occur, in view of the first constraint in (3.2). By a similar argument as before, one finds that, in those two cases, (a) and (b), the sum in (3.15) is thus degenerated into
$\frac{1}{L}$ $\sup_{C\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\ast}\sim C}\frac{S(n,\ell^{2}_{2};c^{\ast}M)}{c^{\ast}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.20)
$\frac{A_{F}(\ell_{1},1)}{\sqrt{L}}$ $\sup_{C\ll\sqrt{\mathcal{X}}L/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c^{\ast}\sim C}\frac{S(n,\ell_{1}\ell^{2}_{2};c^{\ast}M)}{\hat{c}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}/L}\right),$ (3.21)
and
$\frac{1}{L^{2}}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{\hat{c}\sim C}\frac{S(n,1;\hat{c}M)}{\hat{c}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right),$ (3.22)
$\frac{A_{F}(\ell_{1}\ell_{2},1)}{L}$ $\sup_{C\ll\sqrt{\mathcal{X}}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{\hat{c}\sim C}\frac{S(n,\ell_{1}\ell_{2};\hat{c}M)}{\hat{c}M}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}/L^{2}}\right),$ (3.23)
respectively. Recall that, in Case II, the sum (3.15) is explicitly of the following form
$\sup_{C\ll\sqrt{\mathcal{X}}L^{2}/M^{1-\varepsilon}}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\sum_{c\sim C}\frac{S(n,\ell^{2}_{1}\ell^{2}_{2};cM)}{cM}\widetilde{U}_{\ell}\left(\frac{n}{\mathcal{X}}\right).$ (3.24)
On account of the similarities between the expression in (3.24) and that in (3.20)–(3.23), it suffices for us to handle the sum (3.24) with $(c,\ell_{1}\ell_{2})=1$, akin to Case I. In a similar vein, one can deal with the four sub-sums (3.20)–(3.23) in the degenerate case, and the contributions from them are dominated by that of (3.24). Upon combining with this and the argument in Case I as above, our goal eventually is boiled down to estimating (3.15) with $(c,\ell)=1$ in the following stage.
Having seen that the $n$-sum of (3.15) is already in a shape for an application of the the Voronoĭ, upon using (2.9), one might reduce (3.15) essentially to
$\begin{split}\sup_{C\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\frac{CMq}{\sqrt{\mathcal{X}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{mr^{2}}\,e\left(\mp\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)\,\Omega_{\pm}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}};{U}_{\ell}}\right).\end{split}$ (3.25)
Here, we have applied the estimate involving Ramanujan sum for each modulus $c$ that $S(n,0;c)=\sum_{ab=c}\mu(a)\sum_{\beta\bmod c}e{\left({\beta n}/{c}\right)}$. One should notice that the additive twist does not collude with the level anymore, namely, there holds that $(cM,q)=1$; this, however, is guaranteed by the pre-condition in (3.1), which confines that $c<q$ [^3] Indeed, on account of the final restrictions for the parameter $L$ (see (3.46) below), the pre-condition, here, seems dispensable (as far as the establishment of effective ranges for subconvexity is concerned).. By the asymptotic formula, Lemma 2.2, for the integral transform $\Omega_{\pm}$, we are further reduced to the following superiorem problem that
$\begin{split}\mathcal{S}(M,q)\ll\sup_{1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\Psi(C;\ell,M,q),\end{split}$ (3.26)
with
$\Psi(C;\ell,M,q)=\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\,\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}r^{2/3}}\,e\left(\mp\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)\,\widetilde{{U}_{\ell}}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}}}\right).$
Here, $\widetilde{{U}_{\ell}}(x)=\int_{\mathbb{R}^{+}}U_{\ell}(y)e(\pm 3(xy)^{1/3})\mathrm{d}y$ for any $x\in\mathbb{R}^{+}$. By the feature (2.15), it can be expressed roughly as
$\left(\frac{\sqrt{\mathcal{X}\ell}}{CM}\right)^{-1/2}\int_{\mathbb{R}^{+}}U(y)F^{+}_{k}{\left(\frac{4\pi\sqrt{y\mathcal{X}\ell}}{CM}\right)}\,e{\left(\frac{4\pi\sqrt{y\mathcal{X}\ell}}{CM}\pm 3(xy)^{1/3}\right)}\mathrm{d}y$ (3.27)
for any $x\in\mathbb{R}^{+}$, where $F^{+}_{k}(t)=\sqrt{2\pi t}\,\mathscr{W}^{+}_{k}(t)$ for any $t\in\mathbb{R}$, satisfying that $t^{j}{F^{+}}^{(j)}_{k}(t)\ll_{j,k}1$ for any $t\gg 1$ by (2.16).
3.1.1. Non-oscillatory case
In this subsection, we would firstly proceed our argument in the situation where the weight $U_{\ell}(y)$ is of the phase roughly “1”. In other words, we start by working with $C$ localizing at $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$ in (3.26). Our target thus simplifies into
$\begin{split}\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}r^{2/3}}\,e\left(\mp\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)\,\widetilde{{U}_{\ell}}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}}}\right).\end{split}$
One finds that $\widetilde{U}^{(j)}_{\ell}\ll X^{\varepsilon}$ for any $j>0$, by repeated integration by parts many times. It thus follows that one may truncate the variable $m$ at
$m\ll\mathcal{X}^{\varepsilon}+q^{2}(CM)^{3}/\mathcal{X}^{1-\varepsilon}\ll q^{2}\sqrt{\mathcal{X}}\ell^{3/2+\varepsilon},$
at the cost of a negligible error. Our next strategy is to make use of the reciprocity law, followed by employing the GL(3)-Voronoĭ; this would lead to a dual form of (3.15). To be precise, one writes
$e\left(\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)=e\left(-\frac{mr^{2}\overline{cM}}{\ell q^{2}}\right)e\left(\frac{mr^{2}}{cM\ell q^{2}}\right).$
Observe that the harmonic $e\left({mr^{2}}/{cM\ell q^{2}}\right)$ is automatically flat. Indeed, this is not surprising, since that all of the variables in the Kloosterman sums always lie in the “Linnik range”. It thus suffices to evaluate
$\begin{split}\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}r^{2/3}}\,e\left(\pm\frac{mr^{2}\overline{cM}}{\ell q^{2}}\right)\,\widetilde{{U}_{\ell}}\left({\frac{mr^{2}\mathcal{X}}{(cM)^{3}q^{2}}}\right).\end{split}$ (3.28)
For any given $\mathscr{M}\geq 1$, we introduce a smooth function $\eta_{\mathscr{M}}(m)$ such that $\sum_{\mathscr{M}\geq 1}\eta_{\mathscr{M}}(m)=1$, where $\eta_{\mathscr{M}}(m)$ is supported on $m\sim\mathscr{M}$, with $\mathscr{M}$ running through the powers of 2 independently, and satisfies that $\mathscr{M}^{j}\eta_{\mathscr{M}}^{(j)}(n)\ll\mathcal{X}^{\varepsilon}$ for any $j\geq 0$. Thus, if now one defines
${\mathcal{W}_{\pm,r,\ell}}{\left({\rho},\nu\right)}=\eta_{\mathscr{M}}(\rho\mathscr{M})\,\widetilde{U}_{\ell}\left({\frac{\rho\mathscr{M}r^{2}\mathcal{X}}{(\nu CM)^{3}q^{2}}}\right)$
for any ${\rho},\nu\in\mathbb{R}^{+}$, then we can recast (3.28) as
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6+\varepsilon}}{CMq^{1/3}}\,\sup_{\begin{subarray}{c}r\leq CM\\ (r,\ell q)=1\end{subarray}}\frac{1}{r^{2/3}}\,\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\left|\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\right|,$ (3.29)
where
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)=\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,r)}}{m^{1/3}}\,e\left(\pm\frac{mr^{2}\overline{cM}}{\ell q^{2}}\right)\,{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{m}{\mathscr{M}};\frac{c}{C}\right)}$ (3.30)
We will merely consider the $+^{\prime}$ case in the following; the case $-^{\prime}$ can be handled in an entirely analogously manner. Observe that the coprimality relation $(r^{2}\overline{cM},\ell q^{2})=1$ has already held by the prior analysis. We now resort to the Voronoĭ formula in the ramified case, (2.10), which further converts the double sum $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ essentially into
$\begin{split}&\frac{\ell q^{2}}{\mathscr{M}^{1/3}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{n^{\prime}n}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\Omega_{\pm}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};{\mathcal{W}_{\pm,r,\ell}}\right)},\end{split}$
where $\overline{r}$ is such that $r\overline{r}\equiv 1\bmod{\ell q^{2}}$. Via Lemma 2.2, this, however, is controlled by the upper-bound of
$\begin{split}&\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell r}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)},\end{split}$ (3.31)
where the integral $\widetilde{\mathcal{W}_{\pm,r,\ell}}$ is given by
$\widetilde{\mathcal{W}_{\pm,r,\ell}}(v;\iota)=\int_{\mathbb{R}^{+}}{\mathcal{W}_{\pm,r,\ell}}{\left(y;\iota\right)}\,e{\left(\pm 3(vy)^{1/3}\right)}\mathrm{d}y$
for any $v,\iota\in\mathbb{R}^{+}$. Appealing to (3.16), one temporarily sees that $\frac{{\partial^{i}}{\partial^{j}}}{{\partial v^{i}}{\partial\iota^{j}}}\,\widetilde{\mathcal{W}_{\pm,r,\ell}}(v;\iota)\ll_{i,j}\mathcal{X}^{\varepsilon}$ for any $i,j,\varepsilon>0$.
Observe that, unlike other scenarios where one only cares about what happens while the parameter $\mathscr{M}$ varying around its height by directly adding a smooth weight after the Cauchy-Schwarz (see, for example, [43, 20] for relevant details), we need to optimize $\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)$ for any $1\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$ in (3.29). Next, we will proceed the argument by distinguishing two different situations.
When $\mathscr{M}$ is small. We first consider the case where $\mathscr{M}$ is small such that $1\leq\mathscr{M}\leq\mathscr{M}^{\ast}$, say, with $\mathscr{M}^{\ast}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}$. Recall (3.30). We then have
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{7/32}C^{1+\varepsilon}.$ (3.32)
Here, we have used the estimate $|A_{F}(m,n)|\ll(mn)^{7/32+\varepsilon}$, which follows from the best record towards to the generalized Ramanujan conjecture for ${\rm{GL}(2)}$ due to Kim and Sarnak; see [29, Proposition 2].
When $\mathscr{M}$ is large. Next, let us turn to the case where $\mathscr{M}$ is suitably large, with
$\mathscr{M}^{\ast}<\mathscr{M}\ll\frac{q^{2}(CM)^{3}}{r\mathcal{X}^{1-\varepsilon}}.$ (3.33)
We are in a position to rely on the estimates for bilinear Kloosterman sums securing further cancellations in (3.31). It turns out that the quantity we are really concerned about is the following
$\begin{split}&\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll\frac{q^{2}(CM)^{3}}{r\mathcal{X}^{1-\varepsilon}}}\frac{\mathscr{M}^{1/6}}{q\sqrt{\ell r}}\sum_{n^{\prime}|\ell q^{2}r}\,\sum_{c\sim C}\sum_{\pm,n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(r\overline{r^{2}}cM,\pm n;\ell q^{2}r/n^{\prime})\,\widetilde{\mathcal{W}_{\pm,r,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathscr{M}}{\ell^{3}q^{6}r};\frac{c}{C}\right)}.\end{split}$
It is necessary that ${\ell^{3}q^{6}}\gg{(CM)^{3}q^{2}}/\mathcal{X}^{1-10\varepsilon}$, with $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$, so that the sum over $n$ is not vacuous. Now, an application of Lemma 2.6, with $Z=\mathcal{X}^{\varepsilon}$ and $Z_{1}=Z_{2}=1$, reveals that the display above is dominated in the magnitude by
$\ll\sup_{\mathscr{M}^{\ast}\leq\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\sum_{n^{\prime}|\ell q^{2}r}\,\frac{\mathscr{M}^{1/6+\varepsilon}(n^{\prime})^{7/32}}{q\sqrt{\ell r}}\left\{r(\ell q^{2})^{2}\sqrt{\frac{C}{(n^{\prime})^{3}\mathscr{M}}}\right.$
$\phantom{=\;\;}\left.\hskip 150.79968pt+\left(\frac{\ell q^{2}r}{n^{\prime}}\right)^{3/4}(r\overline{r^{2}}M,\ell q^{2}r/n^{\prime})^{1/4}C+\frac{\ell q^{2}r}{n^{\prime}}\right\}$
$\ll\frac{\sqrt{Cr}(\ell q^{2})^{3/2}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
Combining with (3.32), it follows that
$\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll(\mathscr{M}^{\ast})^{2/3}r^{7/32}C^{1+\varepsilon}+\frac{\sqrt{Cr}(\ell q^{2})^{3/2+\varepsilon}}{{\mathscr{M}^{\ast}}^{1/3}}+\frac{\sqrt{Cp}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
Upon choosing
$\mathscr{M}^{\ast}=\frac{(\ell q^{2})^{3/2}}{\sqrt{C}r^{9/32}},$ (3.34)
it turns out that
$\sup_{\mathscr{M}\ll q^{2}(CM)^{3}/r\mathcal{X}^{1-\varepsilon}}\mathcal{K}_{\pm,\ell}(\mathscr{M},C,r)\ll\ell q^{2}C^{2/3+\varepsilon}r^{1/32}+\frac{\sqrt{CM}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right].$
Inserting this expression into (3.29), we obtain
$\sup_{C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}}\,\frac{\mathcal{X}^{1/6}}{CMq^{1/3}}\,\,\sup_{\begin{subarray}{c}r\leq CM\\ (r,\ell q)=1\end{subarray}}\frac{1}{r^{2/3}}\,\left(\ell q^{2}C^{2/3}r^{1/32}+\frac{\sqrt{CM}(qr)^{1/3+\varepsilon}}{\mathcal{X}^{1/6}}\left[q\sqrt{\ell}+(\ell q^{2})^{1/4}C\right]\right).$
Observes that the contribution from the portion for $r$ being large (such that $r\gg\mathcal{X}^{\delta},\delta>0$) is much lesser than that from the dominated case of $r=1$ in the expression above. It suffices to concentrate on this case, and we thus are able to deduce
Lemma 3.1.
For any $\ell$ given as in (3.8), which satisfies that
$\ell\geq\frac{M^{1/4}}{q^{1-\varepsilon}},\quad{\ell^{3}q^{6}}\geq\frac{(CM)^{3}q^{2}}{\mathcal{X}^{1-\varepsilon}},\quad 1\ll\frac{(\ell q^{2})^{3/2}}{\sqrt{C}}\leq\frac{q^{2}(CM)^{3}}{\mathcal{X}^{1+\varepsilon}},$ (3.35)
with $C\asymp\sqrt{\mathcal{X}\ell}/M^{1-\varepsilon}$, we have
$\Psi(C;\ell,M,q)\ll\frac{\ell q^{5/3}\mathcal{X}^{1/6+\varepsilon}}{M}+q^{1+\varepsilon}\sqrt{\frac{\ell}{M}}+\frac{\ell^{2}\mathcal{X}^{1/4+\varepsilon}\sqrt{q}}{M}.$ (3.36)
Here, the third condition of (3.35) stems from the restriction (3.33) in the dominated case of $r=1$, with the choice of $\mathscr{M}^{\ast}$ as shown in (3.34). One notices that the last two conditions in (3.35) requires that
$L\geq\frac{\mathcal{X}^{1/6+\varepsilon}}{q^{4/3}}\quad\text{and}\quad L\geq\frac{\mathcal{X}^{1/10}}{q^{6/5}M^{1/5-\varepsilon}},$ (3.37)
respectively, upon observing that $L^{2}\leq\ell\leq L^{4}$, whenever $\ell=\ell_{1}\ell_{2}$ or $\ell^{2}_{1}\ell^{2}_{2}$, with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$.
3.1.2. Oscillatory case
Next, in this part we will focus on the situation where the parameter $C$ is suitably small such that $C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$ in (3.26). Recall (3.27). In this case, the phase of the weight $U_{\ell}(y)$ will be roughly $\sqrt{\mathcal{X}\ell}/CM\gg\mathcal{X}^{\varepsilon}$. Set $\mathcal{A}:=4\pi\sqrt{\mathcal{X}\ell}/CM$, and put $\mathscr{J}(y)=U(y)F_{k}^{+}(\mathcal{A}\sqrt{y})$ for any $y\in\mathbb{R}^{+}$. To begin with, let us take a glance at the integral
$\int_{\mathbb{R}^{+}}\mathscr{J}(y)e(A\sqrt{y}-3(yx)^{1/3})\mathrm{d}y.$ (3.38)
Notice that $\mathscr{J}$ is compactly supported on $[1/2,5/2]$, and satisfies that $\mathscr{J}^{(j)}\ll_{j,k}\mathcal{X}^{j\varepsilon}$ for any $j>0$. On finds that the integral in $y$ above is negligibly small, unless that $\mathcal{A}\asymp x^{1/3}$, and, in which case it is essentially given by
$x^{-1/6}e(-4x/\mathcal{A}^{2})\mathscr{J}{\left((2x^{1/3}/\mathcal{A})^{6}\right)},$
up to an acceptable error term. It is clear that the integral in (3.38) with the positive sign in the exponential factor is negligibly small, by repeated integration by parts sufficiently many times. Inserting this back into (3.27) gives us the equivalent form of $\Psi(C;\ell,M,q)$ in (3.26) as follows:
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m,r\\ m\geq 1\\ r|cM\end{subarray}}\frac{\overline{A_{F}(m,1)}}{\sqrt{mr^{2}}}\,e\left(\pm\frac{mr^{2}\overline{\ell q^{2}}}{cM}\right)e{\left(-\frac{mr^{2}}{2\pi^{2}cM\ell q^{2}}\right)}\mathscr{J}{\left({\frac{mr^{2}}{\mathcal{M}}}\right)},$ (3.39)
with $\mathcal{M}\asymp\mathcal{X}^{1/2+\varepsilon}\ell^{3/2}q^{2}$ for any $\varepsilon>0$. As hinted by the argument in §3.1.1 above, it suffices to concentrate on the $r=1$ case which, however, is thought to be responsible for the dominated contribution for the display in (3.39). Whilst, for the degenerate case where $r$ is suitably large, our analysis works as well, and the bound that we obtain is dominated by the former (as one will expect). In other words, we are simplified into
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{c\sim C}\sum_{\begin{subarray}{c}m\geq 1\end{subarray}}\frac{\overline{A_{F}(m,1)}}{\sqrt{m}}\,e\left(\frac{m\overline{\ell q^{2}}}{cM}\right)e{\left(-\frac{m}{2\pi^{2}cM\ell q^{2}}\right)}\mathscr{J}{\left({\frac{m}{\mathcal{M}}}\right)}.$
Here, as before, it suffices to consider the $+^{\prime}$ case in (3.39), on account of the resemblance of the treatments of both situations, $+^{\prime}$ and $-^{\prime}$. Recall that one has the coprimality relation $(cM,\ell q)=1$. Now, if one attempts to apply the reciprocity law and then the ${\mathrm{GL}}(3)$-Voronoĭ, we finally arrive at
$\frac{1}{(\mathcal{X}\ell)^{1/4}}\sum_{n^{\prime}|\ell q^{2}}\,\sum_{c\sim C}\sum_{n\geq 1}\frac{A_{F}(n^{\prime},n)}{\sqrt{n}}\,S(cM,\pm n;\ell q^{2}/n^{\prime})\,\widetilde{\mathscr{J}_{\pm,\ell}}{\left(\frac{n{n^{\prime}}^{2}\mathcal{M}}{\ell^{3}q^{6}};\frac{c}{C}\right)},$ (3.40)
akin to (3.31), where the integral transform is given by
$\widetilde{\mathscr{J}_{\pm,\ell}}{\left(x;\nu\right)}=\int_{\mathbb{R}^{+}}\mathscr{J}{\left(y\right)}\,e{\left(\frac{\rho\mathcal{M}y}{\nu CM\ell q^{2}}\pm 3(yx)^{1/3}\right)}\mathrm{d}y$
for some non-zero constant $\rho\in\mathbb{R}$. Observe that the phase of the integral $\widetilde{\mathscr{J}_{\pm,\ell}}$ is still roughly $\sqrt{\mathcal{X}\ell}/CM$, and it is negligibly small, unless that $\mathcal{M}/(CM\ell q^{2})\asymp x^{1/3}$, by repeated integration by parts sufficiently many times. This, in turn, indicates that the sum over $n$ is such that
$n{n^{\prime}}^{2}\asymp\mathscr{N}:=\mathcal{M}^{2}/(CM)^{3}\ll\mathcal{X}^{1+\varepsilon}\ell^{3}q^{4}/(CM)^{3}.$
To proceed, one sees that it is necessary that ${\mathcal{X}\ell^{3}q^{4}}/{(CM)^{3}}\gg 1$ for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$. This implies that $\ell\geq\sqrt{M}/q^{2-\varepsilon}$, which is already ensured by the second pre-condition of (3.2) and the condition we have assumed for the case of $\ell=1$ on Page 17. Recall that $(\ell,q)=1$. Now, returning to (3.40), there are two options for $n^{\prime}$, that is,
  • (1)
    $n^{\prime}=1$;
  • (2)
    $n^{\prime}|\ell,n^{\prime}>1$, with $\ell\geq 2$.
Observe that, here, $n^{\prime}=q\text{ or }q^{2}$ cannot occur, since of that $q^{2}>\mathcal{X}^{1+100\varepsilon}\ell^{3}q^{4}/M^{3}$ by the final choice of $L$ as presented in (3.46) below, upon recalling that $\ell=1,\ell_{1}\ell_{2}$ or $\ell^{2}_{1}\ell^{2}_{2}$ with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, the modulo set defined as in (3.3) before. We will be ready to utilize Lemma 2.7 to the sum in (3.40). It suffices to merely consider Case (1); the contribution from Case (2) would be relatively negligible, upon recalling the remark below the proof of Lemma 2.7. As of now, it is clear that
$x^{i}\nu^{j}\frac{\partial^{i}}{\partial x^{i}}\frac{\partial^{j}}{\partial\nu^{j}}\widetilde{\mathscr{J}_{\pm,\ell}}{\left(x;\nu\right)}\ll(\sqrt{\mathcal{X}\ell}/CM)^{i+j}$
for any $i,j>0$. We now apply (2.13), with $Z=1,Z_{1}=Z_{2}=\sqrt{\mathcal{X}\ell}/CM$, finding that the expression in (3.40) is dominated by
$\ll\frac{1}{(\mathcal{X}\ell)^{1/4-\varepsilon}}\left\{\sqrt{\mathscr{N}\ell q^{2}c}+\left(\frac{\sqrt{\mathcal{X}\ell}}{CM}\right)^{1/2}(\ell q^{2})^{3/4}c+c\sqrt{\mathscr{N}}\right\}$
$\ll\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{7/4}q^{3}}{CM^{3/2}}+\frac{\sqrt{C}(\ell q^{2})^{3/4+\varepsilon}}{\sqrt{M}}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{5/4}q^{2}}{\sqrt{C}M^{3/2}}.$
This is controlled by
$\ll\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{5/4}q^{2}}{M^{3/2}}$
for any $1\leq C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$. In other words, we have established the following
Lemma 3.2.
For any $C\ll\sqrt{\mathcal{X}\ell}/M^{1+\varepsilon}$ with $\ell$ being as in (3.8), we obtain
$\Psi(C;\ell,M,q)\ll\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4+\varepsilon}\ell^{5/4}q^{2}}{M^{3/2}}.$ (3.41)
Altogether, the various bounds in (3.12), (3.26), (3.36) and (3.41) above will result in:
Proposition 3.3.
For any =$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\},\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, with $L$ satisfying that the constraints (3.1), (3.2) and (3.37), we have
$\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f)\ll\mathcal{X}^{\varepsilon}\left(\frac{1}{\sqrt{\ell}}+\frac{\mathcal{X}^{1/6}\ell q^{5/3}}{M}+q\sqrt{\frac{\ell}{M}}+\frac{\ell^{2}\mathcal{X}^{1/4}\sqrt{q}}{M}\right.$
$\phantom{=\;\;}\left.\hskip 59.75095pt+\frac{\mathcal{X}^{1/4}\ell^{7/4}q^{3}}{M^{3/2}}+\frac{\mathcal{X}^{1/4}\ell q^{3/2}}{M}+\frac{\mathcal{X}^{1/4}\ell^{5/4}q^{2}}{M^{3/2}}+\frac{\sqrt{\mathcal{X}}}{M^{3/2}}+\frac{q}{M}\right).$ (3.42)
This shows (1.3) immediately after a little simplification, as required.
3.2. Proof of Corollary 1.3
In this section, we shall turn to the proof of Corollary 1.3. Recalling the definition of the amplifier $\mathfrak{A}_{f}$ as presented in (3.4), which is dependent on $f_{0}\in\mathcal{B}^{\ast}_{k}(M),$ and the property (3.5), via the positivity of the central value $L(1/2,F\otimes f)$ for any $f\in\mathcal{B}^{\ast}_{k}(M)$, we obtain
$\mathfrak{A}_{f_{0}}L(1/2,F\otimes f_{0})\ll\sum_{f\in\mathcal{B}^{\ast}_{k}(M)}\omega^{-1}_{f}\mathfrak{A}_{f}L(1/2,F\otimes f).$
Now, as a direct consequence of Proposition 3.3, by summing over $\ell=1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}$ with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, it turns out that
$L(1/2,F\otimes f)\ll\mathcal{X}^{\varepsilon}\left(\frac{M}{L}+{L^{4}q^{5/3}\mathcal{X}^{1/6}}+qL^{2}\sqrt{{M}}+{L^{8}\mathcal{X}^{1/4}}\sqrt{q}\right.$
$\phantom{=\;\;}\left.\hskip 113.81102pt+\frac{\mathcal{X}^{1/4}L^{7}q^{3}}{\sqrt{M}}+{\mathcal{X}^{1/4}L^{4}q^{3/2}}+\frac{\mathcal{X}^{1/4}L^{5}q^{2}}{\sqrt{M}}+\frac{\sqrt{\mathcal{X}}}{L}+\sqrt{\frac{\mathcal{X}}{M}}+q\right).$ (3.43)
To beat the convexity barrier, the parameter $L$ on the first line on the right-hand side of (3.2) satisfies that
$\begin{split}&L\geq\frac{M^{1+\varepsilon}}{\sqrt{\mathcal{X}}},\quad L\leq\frac{\mathcal{X}^{1/12-\varepsilon}}{q^{5/12}}\\ &L\leq\frac{1}{\sqrt{q}}\left(\frac{\mathcal{X}}{M}\right)^{1/4-\varepsilon},\quad L\leq\left(\frac{\mathcal{X}}{q^{2}}\right)^{1/32-\varepsilon};\end{split}$ (3.44)
whilst, the second line requires that
$\begin{split}&L\leq\frac{\mathcal{X}^{1/28-\varepsilon}M^{1/14}}{q^{3/7}},\quad L\leq\left(\frac{\mathcal{X}}{q^{6}}\right)^{1/16-\varepsilon}\quad L\leq\frac{\mathcal{X}^{1/20-\varepsilon}M^{1/10}}{q^{2/5}}.\end{split}$ (3.45)
Recall that $(M^{3/2}q^{2})^{1-\delta}\leq\mathcal{X}\leq(M^{3/2}q^{2})^{1+\varepsilon}$ for any small constants $\delta,\varepsilon>0$. Collecting all the inequalities (3.1), (3.2), (3.37), (3.44) and (3.45), we infer that the effective range for $L$ should be
$\frac{q^{1+\varepsilon}}{{M^{1/4}}}+\frac{M^{(1+3\delta)/4+\varepsilon}}{q^{1-\delta}}+1\leq L\leq\min{\left\{M^{3/64-\varepsilon},\frac{M^{3/32-\varepsilon}}{{q}^{1/4}},\,\frac{{M}^{1/8}}{q^{5/14+\varepsilon}},\,\frac{M^{7/40-\varepsilon}}{q^{3/10}}\right\}}.$ (3.46)
This yields the range $M^{13/64+\varepsilon}\leq q\leq M^{11/40-\varepsilon}$, and thus gives the requirement as shown in (1.4). The assertion of Corollary 1.3 now follows by combining with (3.2).
4. Proofs of Theorems 1.4 & 1.5
In this section, we shall be dedicated to the proofs of Theorem 1.4 and Theorem 1.5.
4.1. Proof of Theorem 1.4
We shall consider the average of the first moments of $\rm\text{GL}(3)\times GL(2)$ and $\text{GL}(2)$ $L$-functions in a family with a ‘harmonic’ average of $k\sim K$. To be precise, let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Let $f$ be a $\text{GL}(2)$-newform of weight $k\in 2\mathbb{Z}$ and full level. Let $K$ be a sufficiently large parameter. Let $\ell\geq 1$ be an integer such that $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$ for any $\delta,\varepsilon>0$. We attempt to obtain an upper-bound for
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f).$ (4.1)
To begin with, we will give a precise expression for the central value of $L(s,F\otimes f)$. This can be derived in a standard way from [23, Theorem 5.3]. Indeed, analogous to [18, Eqn. (4.11)], one finds that
$L(1/2,F\otimes f)=(1+\varepsilon(F\otimes f))\sum_{m,n\geq 1}\frac{\mu(m)A_{F}(n,1)\lambda_{f}(mn)}{\sqrt{m^{3}n}}V^{\star}\left(\frac{m^{3}n}{k^{3}}\right),$ (4.2)
where
$V^{\star}(y)=\frac{1}{2\pi i}\int_{(3)}\left(\cos\left(\frac{\pi s}{4A}\right)\right)^{-24A}\frac{L_{\infty}(s+1/2,F\otimes f)}{L_{\infty}(1/2,F\otimes f)}L(1+2s,F)y^{-s}\frac{\mathrm{d}s}{s},$
with $A>0$ sufficiently large. Upon shifting the line to $\text{Re}(s)=-1/2$, we get
$V^{\star}(y)=L(1,F)+O(k^{\varepsilon}y^{1/2+\varepsilon})$ (4.3)
for $y\leq 1$. Whilst, for $y>1$, one might shift the line to $\text{Re}(s)=A$, obtaining
$y^{j}{V^{\star}}^{(j)}(y)\ll\left(\frac{k^{3}}{y}\right)^{A}$
for any $j\geq 0$, upon exploiting Stirling’s estimates to (2.5). As described in §2.3, the root number $\varepsilon(F\otimes f)$ is dependent on the weight $k$; thus, by (4.2), we need to turn to seeking an effective control over the following two sums
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.4)
and
$\sum_{k\equiv 0\bmod 2}i^{k}\,W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)\,\sum_{n\geq 1}\frac{A_{F}(n,1)\lambda_{f}(n)}{\sqrt{n}}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.5)
where $k^{3-\delta}\leq\mathcal{Y}\leq k^{3-\varepsilon}$ for arbitrarily small constants $\delta,\varepsilon>0$. By equipping with Petersoon formula, Lemma 2.4, we split the expressions in (4.5) and (4.5) into the diagonal terms and the off-diagonal terms. Via (4.3), both diagonal terms add up to
$\sum_{k\equiv 0\bmod 2}(1+i^{k})\,W{\left(\frac{k-1}{K}\right)}\frac{A_{F}(\ell,1)\lambda_{f}(\ell)}{\sqrt{\ell}}V^{\star}{\left(\frac{\ell}{\mathcal{Y}}\right)}=$ $\frac{L(1,F)A_{F}(\ell,1)\lambda_{f}(\ell)K}{4\sqrt{\ell}}\widehat{W}(0)$
$+O(K^{-1/2+\varepsilon}),$ (4.6)
with $\widehat{W}$ being the Fourier transform of $W$, which is given by $\widehat{W}(x)=\int_{\mathbb{R}}W(y)e(xy)\mathrm{d}y$ for any $x\in\mathbb{R}^{+}$.
We shall merely handle the off-diagonal term for the second sum (4.5) in the following. The same analysis works for the first sum (4.4). The main difference is that, for the first sum, the Lemma 2.8 will be used, which, however, produces a possibly weaker bound since of more error terms thereof; for the second one, we will apply Lemma 2.9 instead. It is verifiable that the off-diagonal term for (4.5) is presented as the following form
$2\pi\sum_{c\geq 1}\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\frac{S(n,\ell;c)}{c}J_{k-1}{\left(\frac{4\pi\sqrt{n\ell}}{c}\right)}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)}.$
Now, one applies Lemma 2.9 as mentioned above, which yields two terms
$\mathcal{T}(K,\ell)=\sum_{c\ll\sqrt{\mathcal{Y}\ell}/K^{2-\varepsilon}}\,\sum_{n\geq 1}\frac{A_{F}(n,1)}{\sqrt{n}}\frac{S(n,\ell;c)}{c}\,e{\left(\frac{2\sqrt{n\ell}}{c}\right)}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)},$ (4.7)
and
$\widehat{\mathcal{T}}(K,\ell)=\sum_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll c\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\,\sum_{n\geq 1}\,\frac{A_{F}(n,1)}{\sqrt{n}}\frac{S(n,\ell;c)}{c}W{\left(\frac{4\pi\sqrt{n\ell}}{cK}\right)}V^{\star}{\left(\frac{n}{\mathcal{Y}}\right)}$ (4.8)
plus two error terms $\text{Err}^{\flat}(K,\ell))$ and $\text{Err}^{\natural}(K,\ell)$, say, which can be estimated as
$\begin{split}\text{Err}^{\flat}(K,\ell))&\ll\frac{1}{K^{2-\varepsilon}}\sum_{n\sim\mathcal{Y}}\frac{|A_{F}(n,1)|}{\sqrt{n}}\sum_{c\sim\sqrt{\mathcal{Y}\ell}/K}\frac{|S(n,\ell;c)|}{c}\ll\frac{\mathcal{Y}^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}},\end{split}$ (4.9)
and
$\begin{split}\text{Err}^{\natural}(K,\ell))&\ll\frac{\sqrt{\ell}}{K^{4-\varepsilon}}\sum_{n\sim\mathcal{Y}}{|A_{F}(n,1)|}\sum_{c\ll\sqrt{\ell}\mathcal{Y}^{1/2+\varepsilon}}\frac{|S(n,\ell;c)|}{c^{2}}\ll\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}},\end{split}$ (4.10)
respectively. Here, $\text{Err}^{\flat}(K,\ell)$ comes from the first error term in (2.19); while, $\text{Err}^{\natural}(K,\ell)$ is for the second one. We have applied bound (2.14) in estimating $\text{Err}^{\natural}(K,\ell)$, when $k$ is sufficiently large.
[…]
${\mathcal{T}}(K,\ell)\ll\frac{1}{\mathcal{Y}^{1/2-\varepsilon}}\left(1+(\mathcal{Y}\ell)^{1/2+\varepsilon}\right)\ll\ell^{1/2+\varepsilon}.$ (4.11)
Next, let us take a look at the term $\widehat{\mathcal{T}}$ in (4.8). By invoking the Voronoĭ summation formula, Lemma 2.1, this sum is thus essentially transformed into
$\begin{split}\sup_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll C\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\frac{C}{\sqrt{\mathcal{Y}}}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast},r\\ m^{\ast}\geq 1\\ r|c\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{m^{\ast}r^{2}}\,e\left(\mp\frac{m^{\ast}r^{2}\overline{\ell}}{c}\right)\,\Omega_{\pm}\left({\frac{m^{\ast}r^{2}\mathcal{Y}}{c^{3}};{V}_{c,\ell}}\right),\end{split}$
with $V_{c,\ell}(x)=V^{\star}(x)W{({4\pi\sqrt{x\mathcal{Y}\ell}}/{(cK)})}$ for any $x\in\mathbb{R}^{+}$. By Lemma 2.2, the display above turns out to be controlled by
$\begin{split}\sup_{\sqrt{\mathcal{Y}\ell}/K^{1-\varepsilon}\ll C\ll\sqrt{\mathcal{Y}\ell}/K^{1-100\varepsilon}}\frac{\mathcal{Y}^{1/6}}{C}&\sum_{c\sim C}\sum_{\begin{subarray}{c}\pm,m^{\ast},r\\ m^{\ast}\geq 1\\ r|c\end{subarray}}\frac{\overline{A_{F}(m^{\ast},r)}}{{m^{\ast}}^{1/3}r^{2/3}}\,e\left(\mp\frac{m^{\ast}r^{2}\overline{\ell}}{c}\right)\,\widehat{V_{\pm,\ell}}\left({\frac{m^{\ast}r^{2}\mathcal{Y}}{c^{3}}}\right),\end{split}$ (4.12)
with the integral transform $\widehat{V_{\pm,\ell}}$ being
$\widehat{V_{\pm,\ell}}(x)=\int_{\mathbb{R}^{+}}V^{\star}(y)W{\left(\frac{4\pi\sqrt{y\mathcal{Y}\ell}}{cK}\right)}\,e{\left(\pm 3(xy)^{1/3}\right)}\mathrm{d}y.$
A direct computation with the help of the Kim–Sarnak bound shows an upper-bound of $O(\mathcal{Y}^{1/6+\varepsilon})$ for (4.12); this, however, is sufficient for our purpose. In summary, upon collecting the estimates in (4.1), (4.9), (4.10) and (4.11), we obtain
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2,F\otimes f)$
$\hskip 65.44142pt\ll\mathcal{Y}^{\varepsilon}\left(\frac{K^{1+\varepsilon}}{\sqrt{\ell}}+\frac{\mathcal{Y}^{3/4}\ell^{1/4}}{K^{5/2-\varepsilon}}+\frac{\mathcal{Y}\sqrt{\ell}}{K^{4-\varepsilon}}+\ell^{1/2+\varepsilon}+\mathcal{Y}^{1/6+\varepsilon}\right).$ (4.13)
Now, after divided by a factor $\sqrt{\ell}$ on both sides of (4.1), we sum over $K^{1-\delta}\ll\ell\ll K^{1+\varepsilon}$ for any $\delta,\varepsilon>0$, obtaining
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2,F\otimes f)\ll K^{1+\varepsilon}.$
This gives (1.6) immediately.
4.2. Proof of Theorem 1.5
At the end of this paper, it remains to complete the proof of Theorem 1.5. This can be realized by a similar argument in §4.1. Actually, by repeating the procedure in evaluating the multiple sum (4.1) with $\ell=1$ as in §4.1, one sees that the sum in (1.7) is boiled down to evaluating $\widehat{\mathcal{T}}(K,1)$, up to the error terms $\text{Err}^{\flat}(K,1)$ and $\text{Err}^{\natural}(K,1)$. We are thus allowed to deduce
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,F\otimes f)=\frac{L(1,F)K}{4}\widehat{W}(0)$ $+\widehat{\mathcal{T}}(K,1)+\text{Err}^{\flat}(K,1)$
$+\text{Err}^{\natural}(K,1)+O(K^{-1/2+\varepsilon}).$
Observe that (4.12) is estimated as $O(\mathcal{Y}^{-100})$ whenever $\ell=1$, which, however, implies that $\widehat{\mathcal{T}}(K,1)$ is negligibly small in term of the magnitude. One concludes that
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,F\otimes f)$ $=\frac{L(1,F)K}{4}\widehat{W}(0)+O(K^{-1/4+\varepsilon}),$
as desired.